### v1

# Semantic Self-Distillation for Language Model Uncertainty

Abstract
Large language models present challenges for principled uncertainty quantification, in part due to their complexity and the diversity of their outputs. Semantic dispersion, or the variance in the meaning of sampled answers, has been proposed as a useful proxy for model uncertainty, but the associated computational cost prohibits its use in latency-critical applications. We show that sampled semantic distributions can be distilled into lightweight student models which estimate a prompt-conditioned uncertainty before the language model generates an answer token. The student model predicts a semantic distribution over possible answers; the entropy of this distribution provides an effective uncertainty signal for hallucination prediction, and the probability density allows candidate answers to be evaluated for reliability. On TriviaQA, our student models match or outperform finite-sample semantic dispersion for hallucination prediction and provide a strong signal for out-of-domain answer detection. We term this technique Semantic Self-Distillation (SSD), which we suggest provides a general framework for distilling predictive uncertainty in complex output spaces beyond language.
1 Introduction
Uncertainty quantification (UQ) can be used to assess the reliability of large language model (LLM) outputs (Xiong et al., 2024; Liu et al., 2025). As LLMs are rapidly deployed in high-stakes domains such as healthcare and law, it becomes increasingly important to have meaningful measures for the trustworthiness of model-generated content. Well-designed UQ methods can provide signals for detecting hallucinations - outputs that are misaligned with fact or user intention - and enable mitigation strategies such as abstention or optimal answer selection (Wen et al., 2025).
LLMs are now frequently embedded in agentic frameworks, which chain multiple model invocations to solve complex, multi-step challenges (Yao et al., 2023; Xi et al., 2025). In these settings a single unreliable output can result in an error which compounds destructively over long-horizon tasks, resulting in complete failure of the system to achieve its goal (Kwa et al., 2025). Given the computational burden of agentic frameworks, practical uncertainty estimators must be low latency compared to the base language models employed.
Semantic dispersion can be a useful proxy for uncertainty: when a language model generates multiple semantically distinct answers to the same prompt, the response is unlikely to be factual and correct. Measures based on semantic dispersion are effective for hallucination detection (Farquhar et al., 2024) but require generating many stochastic completions, rendering them prohibitively slow for latency-sensitive agentic applications.
Probing techniques can address this latency issue by employing lightweight models to predict semantic dispersion or hallucination probability directly from internal model representations (Kadavath et al., 2022; Obeso et al., 2025; Han et al., 2025; Kossen et al., 2024). However, these methods typically collapse the complex notion of language model uncertainty into a single scalar metric, ignoring the topology of the underlying output distribution. This lossy reduction limits their application to simple risk classification, preventing the extraction of richer reliability signals.
In this work, we propose Semantic Self-Distillation (SSD), a general UQ framework which bridges simple probing techniques and expensive sampling-based methods. Instead of regressing a single scalar statistic, SSD trains a small student model to predict a prompt-conditioned distribution over semantic answer embeddings. We model the student as a mixture density network (MDN) (Bishop, 1994), and approximate the teacher distribution using stochastic samples from the base language model. The resulting density acts as a general-purpose uncertainty object: its analytic entropy provides a pre-generation risk forecast, while its posterior likelihood enables post-generation answer verification and other distribution-level operations. Figure 1 provides an overview of the SSD training procedure and the inference-time applications investigated in this work.
On the TriviaQA dataset, SSD provides hallucination-risk signals that in many cases improve upon finite-sample empirical dispersion estimates. Beyond hallucination classification, the distilled density supports (i) context verification via likelihood-based out-of-domain detection and (ii) semantic consensus estimation from a single forward pass.
We summarize our contributions as the following:
  • We introduce Semantic Self-Distillation, a method for distilling a language model’s sampled semantic answer distribution into a lightweight conditional density estimator.
  • We evaluate SSD for pre-generation hallucination prediction, characterizing its cost-accuracy trade-offs relative to sampling-based semantic dispersion and scalar probes.
  • We demonstrate additional reliability primitives enabled by the predicted density, including post-generation verification via likelihood and single-pass semantic consensus estimation.
Figure 1: Overview of Semantic Self-Distillation (SSD). (A) Training: For each prompt, a hidden representation $\mathbf{h}$ conditions a student that distils sampled answer embeddings into a semantic mixture distribution. (B-D) Inference: The predicted density supports pre-generation hallucination prediction via entropy, post-generation context verification via likelihood, and semantic consensus estimation via the mixture mean. Plots illustrate the semantic embedding space, where points denote sampled answers and the mixture mean represents the model’s semantic consensus.
2 Methods
2.1 Problem Setup
Let $x_{1:T}$ be an input prompt and $y_{1:L}$ a completion generated by an autoregressive model. Let $\mathbf{h}=f_{\theta}(x_{1:T})\in\mathbb{R}^{d_{h}}$ denote a vector representation of the prompt extracted from the model’s internal states. We define a semantic embedding $\mathbf{z}\in\mathbb{R}^{d_{z}}$ of the completed sequence via a mapping
$\mathbf{z}=\Phi(y_{1:L}).$
We train a student model $q_{\phi}(\mathbf{z}\mid\mathbf{h})$ to predict the distribution over semantic embeddings $\mathbf{z}$ given only $\mathbf{h}$. The entropy of $q_{\phi}(\cdot\mid\mathbf{h})$ provides an uncertainty signal for hallucination prediction.
2.2 Prompt and Answer Representations
Given a user prompt $x$ of length $T$, we extract hidden states directly from the base language model. Let $\mathbf{h}_{t}^{(\ell)}\in\mathbb{R}^{d}$ denote the hidden state at layer $\ell$ for token position $t$. We utilize the representation of the final token of the prompt extracted from a single specific layer $\ell^{*}$:
$\mathbf{h}\;=\;\mathbf{h}_{T}^{(\ell^{*})}.$
The layer index $\ell^{*}$ is treated as a hyperparameter. In practice, we estimate the optimal $\ell^{*}$ by identifying the layer most predictive of the model’s semantic uncertainty, as detailed in Section 3.
For the answer, we apply a specialist embedding model to the generated sequence to obtain the semantic representation (see Section 3 for model details). To improve training efficiency and focus on the principal axes of semantic variation, we apply Principal Component Analysis (PCA) to the target embeddings $\mathbf{z}$. We fit the PCA transformation on the flattened set of stochastic training completions.
2.3 Student Distribution
We model the student as an MDN (Bishop, 1994). Conditioned on the prompt representation $\mathbf{h}\in\mathbb{R}^{d_{h}}$, the student predicts a $K$-component Gaussian mixture over the semantic embedding $\mathbf{z}\in\mathbb{R}^{d_{z}}$:
$q_{\phi}(\mathbf{z}\mid\mathbf{h})=\sum_{k=1}^{K}\pi_{k}(\mathbf{h})\,\mathcal{N}\!\big(\mathbf{z};\,\bm{\mu}_{k}(\mathbf{h}),\,\operatorname{diag}(\bm{\sigma}_{k}^{2}(\mathbf{h}))\big),$
$\pi_{k}(\mathbf{h})=\mathrm{softmax}_{k}\big(\bm{\alpha}(\mathbf{h})\big),$
where an MLP with parameters $\phi$ maps $\mathbf{h}$ to mixture logits $\bm{\alpha}(\mathbf{h})\in\mathbb{R}^{K}$, component means $\bm{\mu}_{k}(\mathbf{h})\in\mathbb{R}^{d_{z}}$, and (log-)scales $\bm{\sigma}_{k}(\mathbf{h})\in\mathbb{R}^{d_{z}}$. We train by maximizing the conditional log-likelihood of teacher samples. For data efficiency, we use diagonal covariances.
2.4 Teacher Construction and Training
For each training prompt $x$, we extract the representation $\mathbf{h}$ (using the selected layer $\ell^{*}$) and generate $S$ stochastic completions $y^{(1)},\dots,y^{(S)}$. We compute semantic embeddings $\mathbf{z}^{(s)}=\Phi(y^{(s)})$ for $s{=}1,\dots,S$, and fit the student parameters $\phi$ by conditional maximum likelihood:
$\max_{\bm{\phi}}\ \sum_{\mathbf{x}}\ \frac{1}{S}\sum_{s=1}^{S}\log q_{\bm{\phi}}\!\big(\mathbf{z}^{(s)}\mid\mathbf{h}\big).$
2.5 Quantifying Dispersion
Given a new test prompt $x_{1:T}$, we compute its representation $\mathbf{h}=f_{\theta}(x_{1:T})$ in a single forward pass of the base model. The student then predicts the parameters of a $K$-component diagonal-covariance Gaussian mixture over semantic embeddings. We measure dispersion using the order-2 Rényi entropy (Rényi, 1961) of the predictive density:
$\mathsf{H}_{2}\!\big(q_{\phi}(\cdot\mid\mathbf{h})\big)=-\log\int q_{\phi}(\mathbf{z}\mid\mathbf{h})^{2}\,d\mathbf{z},$
Intuitively, the integral $\int q(\mathbf{z})^{2}d\mathbf{z}$ is the collision probability of the density: it is larger when mass is concentrated, and smaller when the distribution is spread out (Principe, 2010). Higher $\mathsf{H}_{2}$ corresponds to greater semantic dispersion.
We use Rényi-2 because it yields an exact closed form for Gaussian mixtures (Wang et al., 2009), enabling a sampling-free, analytic dispersion score at inference time. By contrast, the Shannon differential entropy (Shannon, 1948) does not, in general, admit a closed-form expression for Gaussian mixture models; it is typically approximated using Monte Carlo estimation or via analytic bounds (Huber et al., 2008; Dahlke and Pacheco, 2023). Using Rényi-2 therefore preserves SSD’s goal of avoiding additional sampling or expensive numerical estimation at deployment. We provide the closed form expression and derivation in Appendix Section A.
2.6 Using the Predicted Semantic Distribution
At inference time, SSD produces a prompt-conditioned semantic density $q_{\phi}(\mathbf{z}\mid\mathbf{h})$ over answer embeddings, which we treat as a general-purpose uncertainty object with three primitive operations:
(i) Semantic Dispersion Prediction
We use the Rényi-2 entropy $\mathsf{H}_{2}\!\big(q_{\phi}(\cdot\mid\mathbf{h})\big)$ to provide a prompt-level uncertainty estimate in a single forward pass.
(ii) Context Verification
Given any candidate response $\mathbf{y}^{\prime}$, we embed it as $\mathbf{z}^{\prime}=\Phi(\mathbf{y}^{\prime})$ and score its compatibility with the prompt via the posterior log-likelihood $\log q_{\phi}(\mathbf{z}^{\prime}\mid\mathbf{h})$. We treat answers with low likelihood as out-of-domain relative to the prompt context.
(iii) Latent Sampling
Because $q_{\phi}(\mathbf{z}\mid\mathbf{h})$ is an explicit mixture distribution, we can compute summary statistics such as the mixture mean $\mathbb{E}_{q_{\phi}}[\mathbf{z}]=\sum_{k}\pi_{k}(\mathbf{h})\,\bm{\mu}_{k}(\mathbf{h})$ and, when desired, draw latent samples $\tilde{\mathbf{z}}\sim q_{\phi}(\cdot\mid\mathbf{h})$.
3 Experiments
3.1 Models and Dataset Generation
We evaluate the SSD distribution in the context of short-form question answering, using the TriviaQA dataset (Joshi et al., 2017). We construct a training set of $N_{train}=4000$ prompts and report results on a held-out test set of $N_{test}=1000$ prompts.
We evaluate across five model families: Mistral (Ministral-8B-Instruct), Llama-3 (3.1-8B-Instruct, 3.2-3B-Instruct), Qwen (Qwen3-4B-Instruct-2507, Qwen3-8B), Gemma (Gemma-3-4B) and SmolLM (SmolLM3-3B). For all experiments, we use Qwen3-8B as the oracle ‘judge’ model to determine the correctness of generated answers.
For each prompt and model, we sample $S=32$ samples at temperature $T=1$. We additionally sample a single ‘default’ answer at $T=0.1$ which we input to the judge model to generate the hallucination label.
For the answer embedding, we use EmbeddingGemma (Vera et al., 2025), with the embedding truncated to a dimension of $d_{z}=128$ . For hallucination detection, we further reduce dimensionality using PCA to $d_{z}^{\prime}=16$, following ablation studies detailed in Appendix Section D. For likelihood-based OOD verification, we set $d_{z}^{\prime}=32$.
3.2 Layer Selection Protocol
To estimate a suitable layer $\ell^{*}$ for the student input, we use the results of our hallucination detection baseline experiments, described in Section 3.3. Following Kossen et al. (2024), we first train simple linear probes on all layers to predict the binarized Semantic Entropy. We select the layer index $\ell^{*}$ that maximizes validation accuracy for this proxy task. We assume that the layer most predictive of the scalar entropy value also contains the richest information for constructing the full distributional mixture. We find this typically corresponds to deeper layers, shown in Table 5 in the Appendix.
3.3 Hallucination Detection via Predicted Dispersion
We treat the student’s dispersion score as a pre-generation risk score: higher values indicate greater semantic spread and higher hallucination risk. We report AUROC and AUPRC against correctness labels determined using an LLM-as-judge framework.
SSD compares favourably to teacher dispersion.
We first compare the student’s dispersion score to the standard deviation of the teacher’s $S=32$ semantic embeddings, which we term teacher dispersion (TD). Table 1 reports results across models. Despite requiring no inference-time sampling, SSD outperforms TD on four of the seven models investigated in AUROC and on all seven in AUPRC. We attribute this in part to a smoothing effect: TD is a finite-sample statistic and assigns zero uncertainty to prompts where the model consistently produces the same answer, even when that answer is incorrect. By contrast, SSD learns a continuous mapping from prompt representations to semantic uncertainty, enabling differentiated risk estimates in such degenerate cases, which likely contributes to its stronger AUPRC.
Model AUROC AUPRC
TD SSD TD SSD
Qwen3-8B 0.684 $\pm$ 0.017 0.677 $\pm$ 0.018 0.525 $\pm$ 0.027 0.570 $\pm$ 0.028
Qwen3-4B 0.732 $\pm$ 0.016 0.699 $\pm$ 0.016 0.596 $\pm$ 0.025 0.618 $\pm$ 0.025
Llama 3.1 8B 0.695 $\pm$ 0.017 0.686 $\pm$ 0.020 0.443 $\pm$ 0.028 0.519 $\pm$ 0.030
Llama 3.2 3B 0.665 $\pm$ 0.017 0.738 $\pm$ 0.017 0.558 $\pm$ 0.027 0.657 $\pm$ 0.027
Ministral 8B 0.743 $\pm$ 0.016 0.748 $\pm$ 0.016 0.579 $\pm$ 0.029 0.599 $\pm$ 0.029
SmolLM3 3B 0.698 $\pm$ 0.016 0.711 $\pm$ 0.016 0.569 $\pm$ 0.025 0.623 $\pm$ 0.026
Gemma 3 4B 0.676 $\pm$ 0.017 0.701 $\pm$ 0.017 0.632 $\pm$ 0.024 0.662 $\pm$ 0.025
Table 1: Distilling sampled dispersion into a single pass. TD computes dispersion from $S=32$ sampled semantic embeddings at inference time, while SSD predicts a prompt-conditioned semantic density and scores dispersion analytically. We report hallucination detection AUROC/AUPRC as mean $\pm$ standard deviation over 1,000 bootstrap resamples of the test set; bold indicates the better of TD vs. SSD per metric.
Comparison to additional baselines.
We further compare SSD against three widely used hallucination detection methods:
  1. 1.
    Probability of Correctness Probe (PCP): A logistic regression classifier trained on the prompt representation to predict binary correctness. Originally proposed as ‘P(I know)’ by Kadavath et al. (2022), this method requires externally provided correctness labels.
  2. 2.
    Semantic Entropy (SE): A state-of-the-art sampling-based approach that measures semantic dispersion by clustering $S$ stochastic completions using a natural language inference (NLI) model to assess mutual entailment (Farquhar et al., 2024).
  3. 3.
    Semantic Entropy Probe (SEP): A lightweight probe with the same architecture as PCP, trained to regress a binarized SE target, inheriting SE’s sampling and NLI-based target construction cost (Kossen et al., 2024).
To reflect agentic deployment constraints, we report hallucination detection performance alongside method characteristics: inference latency, supervision target construction cost, and whether a method supports post-generation (posterior) signals in addition to pre-generation (prior) uncertainty. Table 2 summarizes these trade-offs and reports macro-average AUROC across model families; full per-model AUROC/AUPRC tables are provided in the Appendix (tables 8 and 9).
Overall, SE achieves the strongest AUROC, consistent with prior work showing that explicit sampling of multiple reasoning paths yields high-quality uncertainty signals (Kadavath et al., 2022). Among single-pass methods, PCP and SEP outperform SSD as pre-generation hallucination predictors, confirming that the binary classification problem is simpler than distilling the geometry of a full semantic distribution. However, SSD occupies a distinct and practically useful regime: it matches probe-level inference latency, avoids entailment-based target construction, and uniquely provides posterior signals as well as prior risk scores. The post-generation signals enable verification and filtering mechanisms that are not available to scalar probes.
Method AUROC $\uparrow$ Latency $\downarrow$ Target cost $\downarrow$ Posterior
SE 0.804 $\pm$ 0.036 $S$ samples + $S^{2}$ NLI $\times$
SEP 0.759 $\pm$ 0.026 Single-pass $S$ samples + $S^{2}$ NLI $\times$
PCP 0.771 $\pm$ 0.021 Single-pass $C_{\text{CL}}$ $\times$
TD 0.699 $\pm$ 0.027 $S$ samples $\times$
SSD 0.708 $\pm$ 0.024 Single-pass $S$ samples + $S\,\Phi$ $\ell(y),\ \mathbb{E}[\mathbf{z}]$
Table 2: Hallucination detection trade-offs. AUROC is reported as a macro-average (mean $\pm$ std) across models. Latency summarizes inference-time requirements (single-pass vs $S$ sampled generations, with entailment-based SE additionally requiring $S^{2}$ natural language inference (NLI) comparisons). Target cost denotes the computation required to construct supervision targets during training; $C_{CL}$ is the cost of acquiring correctness labels. Only SSD explicitly provides post-generation posterior signals in addition to pre-generation uncertainty.
3.4 Distillation Fidelity vs. Detection Performance
We examine whether the quality of distillation (how closely the student learns the teacher distribution) constrains hallucination detection performance. We define distillation fidelity ($\rho_{\text{fidelity}}$) as the Spearman rank correlation, over $N=1000$ test prompts, between the student’s predicted entropy and the teacher dispersion. A high $\rho_{\text{fidelity}}$ indicates a more faithful recovery of the teacher’s prompt-induced uncertainty ordering.
Figure 2 plots hallucination detection AUROC against distillation fidelity across base models and student hyperparameter settings. We observe a strong positive relationship, with a cross-model Spearman correlation of $\rho_{\text{meta}}=0.65$, indicating that detection performance is limited by the student’s ability to learn the prompt-to-dispersion mapping. The failure cases, notably Qwen3-4B, exhibit near-zero $\rho_{\text{fidelity}}$, suggesting that the relevant uncertainty signal is not accessible from the chosen prompt representation.
Figure 2: Distillation fidelity drives detection performance. The X-axis shows distillation fidelity ($\rho_{\text{fidelity}}$): the Spearman correlation between the student’s predicted entropy and the teacher dispersion across test prompts. Multiple points per model correspond to different student hyperparameter settings. $\rho_{\text{meta}}$ is the Spearman correlation between the X and Y variables; it shows that models where the student fails to learn the distribution (low $\rho_{\text{fidelity}}$) result in poor detection.
3.5 Context Verification via Out-of-Domain Detection
We evaluate the SSD density as a post-generation verification signal by measuring its ability to detect mismatched prompt-answer pairs. We frame this as an out-of-domain detection task: for each test prompt $x_{i}$, we pair its representation $\mathbf{h}_{i}$ with (i) the model’s default answer $y_{i}$ and (ii) a random answer $y_{j}$ drawn from a different prompt. Each candidate answer is embedded as $\mathbf{z}=\Phi(y)$ and scored using the student’s posterior log-likelihood $\log q_{\phi}(\mathbf{z}\mid\mathbf{h})$. We report AUROC for distinguishing matched from mismatched pairs.
As shown in Table 3, the posterior likelihood is a strong context-verification signal, achieving AUROCs above 0.95 for most base models. This suggests that SSD can serve as a low-latency guardrail in agentic workflows, for example detecting cache poisoning (Chen et al., 2024) or filtering retrieved content in retrieval-augmented generation (Lewis et al., 2020).
The likelihood provides a modest signal for post-generation hallucination detection, demonstrated by AUROCs above 0.5. This result could be interpreted through the lens of Semantic Density (Qiu and Miikkulainen, 2024), which posits that reliable samples tend to lie in denser regions of a semantic predictive space. Rather than approximating this distribution via samples and associated sequence likelihoods, our approach explicitly models it.
Table 3: Utility of SSD posterior likelihood. Out-of-domain (OOD) detection reports AUROC for distinguishing an answer belonging to the prompt from a random answer. Hallucination detection compares AUROC for pre-generation entropy vs. post-generation likelihood. While likelihood is a strong OOD detector, entropy remains the superior signal for hallucination classification. Values report mean AUROC with bootstrap standard deviation over 1,000 resamples.
Model OOD Hallucination
Likelihood Entropy Likelihood
Qwen3 8B 0.983 $\pm$ 0.002 0.697 $\pm$ 0.017 0.571 $\pm$ 0.019
Qwen3 4B 0.984 $\pm$ 0.002 0.703 $\pm$ 0.016 0.555 $\pm$ 0.018
Llama 3.1 8B 0.989 $\pm$ 0.002 0.702 $\pm$ 0.019 0.578 $\pm$ 0.020
Llama 3.2 3B 0.980 $\pm$ 0.003 0.728 $\pm$ 0.016 0.612 $\pm$ 0.018
Ministral 8B 0.944 $\pm$ 0.005 0.758 $\pm$ 0.016 0.616 $\pm$ 0.018
SmolLM3 3B 0.988 $\pm$ 0.002 0.676 $\pm$ 0.018 0.563 $\pm$ 0.018
Gemma 3 4B 0.895 $\pm$ 0.007 0.679 $\pm$ 0.017 0.577 $\pm$ 0.018
3.6 Semantic Consensus from the Predicted Mixture
We evaluate the generative utility of the student by measuring its ability to approximate the semantic consensus of the model. We treat the empirical mean of $S=32$ stochastic samples as a reference semantic centroid, then measure the mean squared distance (MSD) in Euclidean space between this centroid and both (i) the model’s low-temperature output and (ii) the analytical mean of the student mixture, $\bm{\mu}_{ssd}=\sum_{k}\pi_{k}\bm{\mu}_{k}$. Lower MSD indicates better recovery of the model’s aggregate semantic belief. MSD is computed as $\mathbb{E}_{i}\left[\|\hat{\mathbf{z}}_{i}-\bm{\mu}^{\text{true}}_{i}\|_{2}^{2}\right]$ over examples $i$ in the subset.
We report results for the best-performing student configuration (Ministral 8B), reflecting the strong architectural dependence of distillation fidelity. As shown in Table 4, the SSD mixture mean provides a closer approximation to the semantic centroid than a single default decoding pass. This effect is most pronounced for incorrect answers, where SSD improves centroid recovery in over 76% of test cases. This aligns with prior work showing that hallucinated outputs tend to deviate from semantic consensus (Lin et al., 2024; Phillips et al., 2025). Together with the OOD results in Table 3, this result suggests that SSD’s pre-generation uncertainty estimates can be complemented by post-generation distance- and likelihood-based checks derived from the same predicted density.
Table 4: Semantic consensus approximation with Ministral 8B. We report the mean squared distance (MSD) of the default answer embedding and the SSD predicted mean to the true semantic centroid (average of $S=32$ samples). Values denote mean MSD, with the bootstrap standard deviation expressed as a percentage of the mean shown as a subscript.
Subset Default MSD $\downarrow$ SSD MSD $\downarrow$ Imp. (%) $\uparrow$ SSD Win Rate (%) $\uparrow$
All 0.0153.2 0.0081.9 +45.5 58.3
Correct 0.0134.4 0.0092.3 +32.3 48.8
Incorrect 0.0184.2 0.0073.4 +63.0 76.5
4 Future Work
Improving distillation fidelity. Our results indicate that SSD’s utility as a pre-generation hallucination signal is primarily limited by distillation fidelity: detection performance improves when the student accurately recovers the teacher’s prompt-induced uncertainty ordering. Promising directions include scaling the distillation dataset, enriching the student input beyond a single-token/single-layer representation, and distilling from stronger teachers (larger $S$, or higher-quality semantic targets). These extensions would clarify whether observed failures reflect an information bottleneck in the prompt representation or insufficient supervision for density estimation.
Combining prior and posterior signals. SSD exposes complementary reliability signals from a single predicted density: pre-generation dispersion forecasts risk, while post-generation likelihood and distance-to-consensus verify prompt–answer agreement. A natural next step is to integrate these signals into decision policies for agents, for example abstaining or retrieving evidence under high dispersion, and accepting or rejecting candidate actions based on likelihood or consensus distance.
Diffusion models. While we focus on autoregressive models, SSD may be particularly well suited to diffusion-based language models, where hidden prompt representations may more directly parameterize the final semantic output distribution (Sahoo et al., 2024).
Domain-specific uncertainty estimation. The technique naturally extends to domain-specific foundation models, such as those in healthcare (Renc et al., 2024; Steinberg et al., 2024) and time-series forecasting (Faw et al., 2025). By distilling predictive distributions over latent outcome representations, students could provide real-time uncertainty estimates for future patient trajectories or other temporal processes, enabling risk-aware decision-making in sequential settings (Zeng et al., 2025).
5 Conclusion
We introduced Semantic Self-Distillation (SSD), a distributional probing framework that distils a language model’s sampled semantic answer distribution into a lightweight, prompt-conditioned density estimator. Unlike scalar probes which compress uncertainty into a single score, SSD predicts an explicit semantic density, enabling analytic dispersion scoring and additional distribution-level operations in a single forward pass.
Empirically, SSD provides competitive hallucination-risk signals relative to finite-sample baselines on several model families, while avoiding the inference-time overhead of generating multiple completions. Crucially, modelling the full density unlocks capabilities that scalar probes cannot match: the posterior likelihood provides strong context verification checks for out-of-domain answers, while mixture statistics allow for single-pass semantic consensus estimates.
Overall, SSD demonstrates that distilling a semantic predictive distribution can yield a general-purpose uncertainty object with applications in forecasting and verification. By decoupling these capabilities from expensive inference-time sampling, SSD offers a practical framework for integrating principled uncertainty quantification into latency-critical agentic applications.
6 Acknowledgments
DAC was funded by an NIHR Research Professorship; a Royal Academy of Engineering Research Chair; and the InnoHK Hong Kong Centre for Cerebro-cardiovascular Engineering (COCHE); and was supported by the National Institute for Health Research (NIHR) Oxford Biomedical Research Centre (BRC) and the Pandemic Sciences Institute at the University of Oxford. EP was funded by an NIHR Research Studentship. SW was supported by the Rhodes Scholarship.
Appendix A Derivations
Closed-form computation of Rényi–2 entropy.
For a Gaussian mixture, the quadratic functional $\int q(\mathbf{z})^{2}d\mathbf{z}$ decomposes into pairwise overlaps between mixture components:
$\int q_{\phi}(\mathbf{z}\mid\mathbf{h})^{2}\,d\mathbf{z}=\sum_{i=1}^{K}\sum_{j=1}^{K}\pi_{i}(\mathbf{h})\,\pi_{j}(\mathbf{h})\int\mathcal{N}(\mathbf{z};\bm{\mu}_{i},\bm{\Sigma}_{i})\,\mathcal{N}(\mathbf{z};\bm{\mu}_{j},\bm{\Sigma}_{j})\,d\mathbf{z}.$
Using the standard Gaussian product identity,
$\int\mathcal{N}(\mathbf{z};\bm{\mu}_{i},\bm{\Sigma}_{i})\,\mathcal{N}(\mathbf{z};\bm{\mu}_{j},\bm{\Sigma}_{j})\,d\mathbf{z}=\mathcal{N}\!\big(\bm{\mu}_{i};\bm{\mu}_{j},\bm{\Sigma}_{i}+\bm{\Sigma}_{j}\big),$
we obtain:
$\mathsf{H}_{2}\!\big(q_{\phi}(\cdot\mid\mathbf{h})\big)=-\log\!\left(\sum_{i=1}^{K}\sum_{j=1}^{K}\pi_{i}(\mathbf{h})\,\pi_{j}(\mathbf{h})\,\mathcal{N}\!\big(\bm{\mu}_{i}(\mathbf{h});\bm{\mu}_{j}(\mathbf{h}),\bm{\Sigma}_{i}(\mathbf{h})+\bm{\Sigma}_{j}(\mathbf{h})\big)\right).$
Equivalently, defining the collision matrix $\mathbf{K}(\mathbf{h})\in\mathbb{R}^{K\times K}$ with entries
$K_{ij}(\mathbf{h})=\mathcal{N}\!\big(\bm{\mu}_{i}(\mathbf{h});\bm{\mu}_{j}(\mathbf{h}),\bm{\Sigma}_{i}(\mathbf{h})+\bm{\Sigma}_{j}(\mathbf{h})\big),$
and $\bm{\pi}(\mathbf{h})\in\mathbb{R}^{K}$ the mixture weights, we have
$\mathsf{H}_{2}\!\big(q_{\phi}(\cdot\mid\mathbf{h})\big)=-\log\!\big(\bm{\pi}(\mathbf{h})^{\!\top}\mathbf{K}(\mathbf{h})\,\bm{\pi}(\mathbf{h})\big).$
This computation is $\mathcal{O}(K^{2})$ in the number of mixture components and is negligible relative to the transformer backbone.
Appendix B Model Metadata
Table 5: Model characteristics. ‘Base Acc’ denotes QA accuracy. ‘PCP’ and ‘SEP’ indicate optimal probe layers. The ‘Best SSD Config’ columns show the student architecture (HH=Hidden, DD=Depth, KK=Components) that achieved the highest SSD AUROC.
Probe Layers Best SSD Config
Model Base Acc PCP SEP H D K AUROC
Qwen3-8B 0.620 32 24 128 6 10 0.677
Qwen3-4B-Instruct-2507 0.565 18 20 256 2 10 0.699
Llama 3.1 8B 0.700 20 16 256 4 2 0.687
Llama 3.2 3B 0.585 15 19 512 3 5 0.737
Ministral 8B 0.656 33 34 1024 6 1 0.748
SmolLM3 3B 0.567 20 19 512 6 5 0.710
Gemma 3 4B 0.532 22 19 512 2 10 0.700
Appendix C Computational Complexity Analysis
The primary motivation for Semantic Self-Distillation is to enable generic uncertainty estimation in AR models without sampling. Here we compare the asymptotic complexity of our method against sampling-based baselines. Let $T_{\text{pre}}$ be the prompt length, $T_{\text{gen}}$ be the generation length, $S$ be the number of samples, and $C_{\text{LLM}}$ be the cost of a single transformer forward pass per token.
Sampling-based Methods (SE, TD).
Computing Semantic Entropy (Farquhar et al., 2024) requires $S$ stochastic sequences per prompt at inference time (typically $S\in[5,20]$). The cost is dominated by the autoregressive decoding steps:
$\mathcal{O}_{\text{SE}}\approx S\times(T_{\text{pre}}+T_{\text{gen}})\times C_{\text{LLM}}+\mathcal{O}_{\text{NLI}}(S^{2})$
where $\mathcal{O}_{\text{NLI}}$ represents the quadratic cost of comparing all sample pairs using an entailment model. This linear scaling with $S$ introduces latency that is often unacceptable for interactive agents.
Scalar Probes (PCP, SEP).
Scalar probes operate on prompt representations extracted from a single forward pass of the base model. The probe itself is a lightweight classifier or regressor applied to a cached hidden state, without requiring additional sampling or decoding:
$\mathcal{O}_{\text{SEP}}\approx 1\times T_{\text{pre}}\times C_{\text{LLM}}+C_{\text{MLP}}$
While efficient, these methods predict a compressed scalar statistic and discard the geometric structure of the model’s semantic output distribution.
Semantic Self-Distillation (SSD).
SSD matches the inference complexity of scalar probes while retaining the distributional richness of sampling methods. The MDN head introduces a slightly larger parameter set than a linear probe, but this cost ($C_{\text{MDN}}$) is negligible compared to the transformer backbone. Crucially, the entropy calculation is analytical and does not require sampling:
$\mathcal{O}_{\text{SSD}}\approx 1\times T_{\text{pre}}\times C_{\text{LLM}}+C_{\text{MDN}}$
Thus, SSD effectively moves the computational burden from inference time (where latency matters) to training time, enabling $O(1)$ uncertainty estimation during deployment.
Appendix D Ablations and Architecture Analysis
D.1 PCA Dimension
Standard sentence embeddings are high-dimensional (e.g., $d_{z}=768$). Training an MDN to model a full covariance density in this space is data-intensive. We apply PCA to reduce the target space to $d_{pca}\in\{16,32,64,128\}$.
This introduces a trade-off: higher dimensions preserve more semantic detail (increasing the theoretical upper bound of detection performance, measured by teacher dispersion), but make the regression task significantly harder for the student. As shown in Figure 3 and Table 6, while the detection performance of teacher dispersion improves monotonically with dimensionality, the student’s performance (SSD) peaks at $d_{pca}=16$. Beyond this point, the student fails to effectively distil the increasingly sparse density, leading to a drop in detection accuracy. For the OOD experiments we find PCA dimensions of 16 and 32 perform similarly, followed by a drop-off at higher dimensions. We fix $d_{pca}=16$ for all main experiments bar the OOD task, where we set $d_{pca}=32$.
We underline that the same training dataset size ($N_{train}=4000$) was used in all PCA ablations; the learning task is therefore more difficult in higher dimensions, as models must learn more complex relationships using the same number of datapoints. For larger training datasets, we expect the SSD performance will continue to match or exceed the TD performance.
Table 6: Effect of PCA dimensionality on Hallucination Detection (SSD) and OOD Verification. Scores are averaged across the best student configuration for each model.
PCA Dim Teacher Dispersion (TD) Peak SSD AUROC Peak OOD AUROC
16 0.699 $\pm$ 0.029 0.708 $\pm$ 0.026 0.965 $\pm$ 0.026
32 0.724 $\pm$ 0.027 0.706 $\pm$ 0.029 0.966 $\pm$ 0.035
64 0.747 $\pm$ 0.025 0.701 $\pm$ 0.037 0.960 $\pm$ 0.047
128 0.757 $\pm$ 0.025 0.703 $\pm$ 0.029 0.950 $\pm$ 0.055
Figure 3: As PCA dimension increases, the semantic resolution improves and the teacher dispersion baseline rises. The distillation task however becomes harder, causing SSD performance to fall.
D.2 Student Capacity and Information Saturation
We investigate whether the performance gap on difficult models can be closed by increasing the student’s capacity. We analyze two axes of complexity:
  1. 1.
    Distributional Complexity: We vary the number of mixture components $K\in\{1,2,5,10\}$.
  2. 2.
    Backbone Capacity: We vary the MLP hidden dimension $H\in\{128,256,512,1024\}$ and depth $D\in\{2,3,4,6\}$.
Table 5 reveals that the optimal hyperparameters vary by base model. Aggregating results across all models (Table 7) reveals capacity saturation, where increasing the network width or depth yields negligible gains in mean AUROC. This saturation suggests that the bottleneck is not the student’s expressivity, but the mutual information available in the final-token prompt representation. The student appears to have extracted most of the readily accessible signal; further improvements would require richer inputs (e.g., attention-pooled representations) rather than larger students. Although we observe optimal performance for $K=10$ mixture components, we opt not to increase this further given our set number of $S=32$ samples; modelling more than ten unique semantic concepts in this setting is unrealistic and would likely damage the generalization ability of the trained student.
Table 7: Marginal impact of student capacity (Fixed PCA=16). Scores represent the mean AUROC across all runs with the specified parameter; standard deviations indicate variability across different base models and remaining hyperparameter configurations. We observe negligible performance gains from increasing complexity, indicating information saturation in the prompt representation.
(a) Components ($K$)
$K$ AUROC
1 0.648 $\pm$ 0.050
2 0.657 $\pm$ 0.052
5 0.660 $\pm$ 0.048
10 0.666 $\pm$ 0.042
(b) Width ($H$)
$H$ AUROC
128 0.662 $\pm$ 0.043
256 0.660 $\pm$ 0.046
512 0.660 $\pm$ 0.050
1024 0.649 $\pm$ 0.054
(c) Depth ($D$)
$L$ AUROC
2 0.657 $\pm$ 0.049
3 0.660 $\pm$ 0.047
4 0.658 $\pm$ 0.050
6 0.656 $\pm$ 0.048
Figure 4: Student capacity ablation, averaged across all seven models investigated.
Appendix E Full Hallucination Detection Results
Table 8: Hallucination Detection AUROC. Values report mean AUROC (%) over 1,000 bootstrap resamples of the test set, with the bootstrap standard deviation shown as a subscript. We mark the best overall result in bold and the best single-pass result with an underline. SSD achieves competitive performance with teacher dispersion but lags scalar probes and SE.
Model PCP SEP SSD (Ours) TD SE
Qwen3-8B 76.91.5 76.21.5 67.71.8 68.41.7 81.61.4
Qwen3-4B 80.21.4 78.31.5 69.91.6 73.21.6 78.51.5
Llama 3.1 8B 77.01.6 74.51.7 68.62.0 69.51.7 85.21.3
Llama 3.2 3B 76.71.5 75.71.5 73.81.7 66.51.7 80.71.4
Ministral 8B 74.01.7 77.11.5 74.81.6 74.31.6 80.51.5
SmolLM3 3B 79.71.4 78.91.4 71.11.6 69.81.6 83.21.2
Gemma 3 4B 75.41.5 70.71.7 70.11.7 67.61.7 73.11.6
Table 9: Hallucination Detection AUPRC. Values report mean AUPRC (%) over 1,000 bootstrap resamples of the test set, with the bootstrap standard deviation shown as a subscript. We mark the best overall result in bold and the best single-pass result with an underline. SSD consistently outperforms the teacher dispersion (TD) baseline, effectively shifting the cost of dispersion estimation from inference time to training time.
Model PCP SEP SSD (Ours) TD SE
Qwen3-8B 64.92.7 64.92.7 57.02.8 52.52.7 75.02.1
Qwen3-4B 74.32.3 70.82.5 61.82.5 59.62.5 73.32.1
Llama 3.1 8B 59.63.1 54.03.0 51.93.0 44.32.8 73.52.7
Llama 3.2 3B 68.92.4 65.42.6 65.72.7 55.82.7 74.22.4
Ministral 8B 59.22.9 62.92.8 59.92.9 57.92.9 69.62.7
SmolLM3 3B 75.12.2 73.32.2 62.32.6 56.92.5 78.91.9
Gemma 3 4B 69.42.4 65.02.5 66.22.5 63.22.4 69.22.3

### v2

@@
 Semantic dispersion, or the variance in the meaning of sampled answers, has been proposed as a useful proxy for model uncertainty, but the associated computational cost prohibits its use in latency-critical applications.
-We show that sampled semantic distributions can be distilled into lightweight student models which estimate a prompt-conditioned uncertainty before the language model generates an answer token.
-The student model predicts a semantic distribution over possible answers; the entropy of this distribution provides an effective uncertainty signal for hallucination prediction, and the probability density allows candidate answers to be evaluated for reliability.
-On TriviaQA, our student models match or outperform finite-sample semantic dispersion for hallucination prediction and provide a strong signal for out-of-domain answer detection.
-We term this technique Semantic Self-Distillation (SSD), which we suggest provides a general framework for distilling predictive uncertainty in complex output spaces beyond language.
+We show that sampled semantic distributions can be distilled into lightweight student models which estimate a prompt-conditioned density before the language model generates an answer token.
+The student model predicts a semantic distribution over possible answers; the entropy of this distribution provides a prompt-level uncertainty signal, and the probability density allows answer-level reliability evaluation.
+Across experiments on TriviaQA and MMLU, we find our student models perform competitively relative to sampling-based semantic dispersion baselines on a hallucination prediction task, whilst offering additional uncertainty primitives for out-of-domain detection and multiple-choice answer selection.
+We term this technique Semantic Self-Distillation (SSD), which can serve as a general framework for distilling predictive uncertainty in complex output spaces beyond language.
 1 Introduction
-Uncertainty quantification (UQ) can be used to assess the reliability of large language model (LLM) outputs (Xiong et al., 2024; Liu et al., 2025).
-As LLMs are rapidly deployed in high-stakes domains such as healthcare and law, it becomes increasingly important to have meaningful measures for the trustworthiness of model-generated content.
-Well-designed UQ methods can provide signals for detecting hallucinations - outputs that are misaligned with fact or user intention - and enable mitigation strategies such as abstention or optimal answer selection (Wen et al., 2025).
-LLMs are now frequently embedded in agentic frameworks, which chain multiple model invocations to solve complex, multi-step challenges (Yao et al., 2023; Xi et al., 2025).
-In these settings a single unreliable output can result in an error which compounds destructively over long-horizon tasks, resulting in complete failure of the system to achieve its goal (Kwa et al., 2025).
-Given the computational burden of agentic frameworks, practical uncertainty estimators must be low latency compared to the base language models employed.
-Semantic dispersion can be a useful proxy for uncertainty: when a language model generates multiple semantically distinct answers to the same prompt, the response is unlikely to be factual and correct.
-Measures based on semantic dispersion are effective for hallucination detection (Farquhar et al., 2024) but require generating many stochastic completions, rendering them prohibitively slow for latency-sensitive agentic applications.
-Probing techniques can address this latency issue by employing lightweight models to predict semantic dispersion or hallucination probability directly from internal model representations (Kadavath et al., 2022; Obeso et al., 2025; Han et al., 2025; Kossen et al., 2024).
-However, these methods typically collapse the complex notion of language model uncertainty into a single scalar metric, ignoring the topology of the underlying output distribution.
-This lossy reduction limits their application to simple risk classification, preventing the extraction of richer reliability signals.
-In this work, we propose Semantic Self-Distillation (SSD), a general UQ framework which bridges simple probing techniques and expensive sampling-based methods.
-Instead of regressing a single scalar statistic, SSD trains a small student model to predict a prompt-conditioned distribution over semantic answer embeddings.
-We model the student as a mixture density network (MDN) (Bishop, 1994), and approximate the teacher distribution using stochastic samples from the base language model.
-The resulting density acts as a general-purpose uncertainty object: its analytic entropy provides a pre-generation risk forecast, while its posterior likelihood enables post-generation answer verification and other distribution-level operations.
-Figure 1 provides an overview of the SSD training procedure and the inference-time applications investigated in this work.
-On the TriviaQA dataset, SSD provides hallucination-risk signals that in many cases improve upon finite-sample empirical dispersion estimates.
-Beyond hallucination classification, the distilled density supports (i) context verification via likelihood-based out-of-domain detection and (ii) semantic consensus estimation from a single forward pass.
-We summarize our contributions as the following:
-  • We introduce Semantic Self-Distillation, a method for distilling a language model’s sampled semantic answer distribution into a lightweight conditional density estimator.
-  • We evaluate SSD for pre-generation hallucination prediction, characterizing its cost-accuracy trade-offs relative to sampling-based semantic dispersion and scalar probes.
-  • We demonstrate additional reliability primitives enabled by the predicted density, including post-generation verification via likelihood and single-pass semantic consensus estimation.
 Figure 1: Overview of Semantic Self-Distillation (SSD).
 (A) Training: For each prompt, a hidden representation $\mathbf{h}$ conditions a student that distils sampled answer embeddings into a semantic mixture distribution.
-(B-D) Inference: The predicted density supports pre-generation hallucination prediction via entropy, post-generation context verification via likelihood, and semantic consensus estimation via the mixture mean.
-Plots illustrate the semantic embedding space, where points denote sampled answers and the mixture mean represents the model’s semantic consensus.
+(B-D) Inference: The predicted density supports pre-generation hallucination prediction via entropy, post-generation context verification via likelihood, and multiple-choice answer selection.
+Plots illustrate the semantic embedding space, where points denote sampled answers.
+Uncertainty quantification (UQ) can be used to assess the reliability of large language model (LLM) outputs [xiong2024efficient, liullmuncertainty2025].
+As LLMs are rapidly deployed in high-stakes domains such as healthcare and law, it becomes increasingly important to have meaningful measures for the trustworthiness of model-generated content.
+Well-designed UQ methods can provide signals for detecting errors like hallucinations - outputs that are misaligned with fact or user intention - and enable mitigation strategies such as abstention, information retrieval, or answer selection [wen-etal-2025-know].
+LLMs are now frequently embedded in agentic frameworks, which chain multiple model invocations to solve complex, multi-step tasks [yao2023react, xi2025rise].
+In these settings a single unreliable output can introduce an error that compounds destructively over long-horizon tasks, potentially resulting in complete failure of the system to achieve its goal [kwa2025measuring].
+Given the computational burden of such agentic frameworks, practical uncertainty estimators must be low latency compared to the base language models.
+Semantic dispersion has emerged as a useful proxy for uncertainty: when a language model generates semantically diverse answers to the same prompt, the response is unlikely to be factual and correct.
+Sampling-based measures of semantic dispersion are effective for hallucination detection [farquhar_detecting_2024], but require generating multiple stochastic completions, rendering them prohibitively expensive in latency-sensitive applications.
+Probing techniques can address this latency issue by training lightweight models to predict semantic dispersion or hallucination probability directly from internal model representations [kadavath2022languagemodelsmostlyknow, obeso2025realtimedetectionhallucinatedentities, han-etal-2025-simple, kossen2024semanticentropyprobesrobust].
+However, these methods typically collapse the complex notion of language model uncertainty into a single scalar metric, ignoring the topology of the underlying output distribution.
+This information reduction limits their application to simple risk classification, preventing the extraction of richer reliability signals.
+In this work, we propose Semantic Self-Distillation (SSD), a general UQ framework that bridges lightweight probing techniques and expensive sampling-based methods.
+Instead of regressing a scalar uncertainty statistic, SSD trains a small student model to predict a prompt-conditioned distribution over semantic answer embeddings.
+We model the student as a mixture density network (MDN) [bishop1994mixture], and approximate the teacher distribution using stochastic samples from the base language model.
+The predicted density serves as a general-purpose uncertainty object: its analytic entropy provides a pre-generation risk forecast, whilst estimated likelihoods enable post-generation answer verification.
+Figure 1 provides an overview of the SSD training procedure and the inference-time applications investigated in this work.
+We evaluate SSD on a hallucination prediction task with TriviaQA and MMLU across multiple LLM families, demonstrating competitive, and in some cases even superior, performance relative to sampling-based semantic dispersion baselines.
+While specialized scalar probes can be strong pre-generation classifiers, SSD uniquely retains distributional structure in a single forward pass.
+This enables additional post-generation capabilities unavailable to standard single-pass methods, which we demonstrate by evaluating SSD on likelihood-based out-of-domain answer detection and multiple-choice answer selection.
+We summarize our contributions as the following:
+  • We introduce Semantic Self-Distillation (SSD), a method for distilling a language model’s sampled semantic answer distribution into a lightweight, prompt-conditioned density estimator.
+  • We show that SSD provides competitive pre-generation hallucination prediction performance across multiple model families, while matching the inference latency of single-pass probing methods.
+  • We demonstrate that modelling the full semantic distribution unlocks new reliability primitives beyond scalar uncertainty scoring, including post-generation answer verification and distribution-based answer selection.
 2 Methods
+In this section, we describe how SSD distils a language model’s sampled semantic output distribution into a prompt-conditioned density estimator, rather than a scalar uncertainty proxy.
+We first define the representations used for prompts and answers, specify the MDN student model, and detail the distillation objective used for training.
+We then show how the resulting predictive density can be used to extract both pre-generation, prompt-level dispersion estimates, and post-generation, answer-level likelihood scores, all in a single forward pass at inference time.
 2.1 Problem Setup
-Let $x_{1:T}$ be an input prompt and $y_{1:L}$ a completion generated by an autoregressive model.
+Let $x_{1:T}$ be an input prompt and $y_{1:L}$ a completion generated by a decoding model.
 Let $\mathbf{h}=f_{\theta}(x_{1:T})\in\mathbb{R}^{d_{h}}$ denote a vector representation of the prompt extracted from the model’s internal states.
-We define a semantic embedding $\mathbf{z}\in\mathbb{R}^{d_{z}}$ of the completed sequence via a mapping
+We define a semantic representation $\mathbf{z}\in\mathbb{R}^{d_{z}}$ of the completed sequence via a mapping $\Phi$, which in practice is an off-the-shelf sequence embedding model:
 $\mathbf{z}=\Phi(y_{1:L}).$
 We train a student model $q_{\phi}(\mathbf{z}\mid\mathbf{h})$ to predict the distribution over semantic embeddings $\mathbf{z}$ given only $\mathbf{h}$.
-The entropy of $q_{\phi}(\cdot\mid\mathbf{h})$ provides an uncertainty signal for hallucination prediction.
-2.2 Prompt and Answer Representations
-Given a user prompt $x$ of length $T$, we extract hidden states directly from the base language model.
+2.1.1 Prompt and Answer Representations
+Given a user prompt $x_{1:T}$ of length $T$, we extract hidden states directly from the base language model.
 Let $\mathbf{h}_{t}^{(\ell)}\in\mathbb{R}^{d}$ denote the hidden state at layer $\ell$ for token position $t$.
@@
 The layer index $\ell^{*}$ is treated as a hyperparameter.
-In practice, we estimate the optimal $\ell^{*}$ by identifying the layer most predictive of the model’s semantic uncertainty, as detailed in Section 3.
-For the answer, we apply a specialist embedding model to the generated sequence to obtain the semantic representation (see Section 3 for model details).
-To improve training efficiency and focus on the principal axes of semantic variation, we apply Principal Component Analysis (PCA) to the target embeddings $\mathbf{z}$.
-We fit the PCA transformation on the flattened set of stochastic training completions.
-2.3 Student Distribution
-We model the student as an MDN (Bishop, 1994).
+In practice, we estimate the optimal $\ell^{*}$ by identifying the layer most predictive of the model’s semantic uncertainty, as explained in Appendix D.1.
+For the answer, we apply a specialist embedding model $\Phi$ to the generated sequence $y_{1:L}$ to obtain the semantic representation, $\mathbf{z}=\Phi(y_{1:L})$, with full model details provided in Section 4.1.
+To improve training efficiency and focus on the principal axes of semantic variation, we also apply Principal Component Analysis (PCA) to reduce the dimensionality of the target embeddings $\mathbf{z}$.
+2.1.2 Student Model
+We define the student model $q_{\phi}(\mathbf{z}\mid\mathbf{h})$ using an MDN [bishop1994mixture].
 Conditioned on the prompt representation $\mathbf{h}\in\mathbb{R}^{d_{h}}$, the student predicts a $K$-component Gaussian mixture over the semantic embedding $\mathbf{z}\in\mathbb{R}^{d_{z}}$:
@@
 $\pi_{k}(\mathbf{h})=\mathrm{softmax}_{k}\big(\bm{\alpha}(\mathbf{h})\big),$
-where an MLP with parameters $\phi$ maps $\mathbf{h}$ to mixture logits $\bm{\alpha}(\mathbf{h})\in\mathbb{R}^{K}$, component means $\bm{\mu}_{k}(\mathbf{h})\in\mathbb{R}^{d_{z}}$, and (log-)scales $\bm{\sigma}_{k}(\mathbf{h})\in\mathbb{R}^{d_{z}}$.
-We train by maximizing the conditional log-likelihood of teacher samples.
-For data efficiency, we use diagonal covariances.
-2.4 Teacher Construction and Training
-For each training prompt $x$, we extract the representation $\mathbf{h}$ (using the selected layer $\ell^{*}$) and generate $S$ stochastic completions $y^{(1)},\dots,y^{(S)}$.
+where an MLP with parameters $\phi$ maps $\mathbf{h}$ to mixture logits $\bm{\alpha}(\mathbf{h})\in\mathbb{R}^{K}$, component means $\bm{\mu}_{k}(\mathbf{h})\in\mathbb{R}^{d_{z}}$, and scales $\bm{\sigma}_{k}(\mathbf{h})\in\mathbb{R}^{d_{z}}$.
+For data efficiency, we use diagonal covariance matrices.
+For each training prompt $x$, we extract the representation $\mathbf{h}$ (using the selected layer $\ell^{*}$) and generate $S$ stochastic completions $y^{(1)},\dots,y^{(S)}$ using the base language model.
 We compute semantic embeddings $\mathbf{z}^{(s)}=\Phi(y^{(s)})$ for $s{=}1,\dots,S$, and fit the student parameters $\phi$ by conditional maximum likelihood:
-$\max_{\bm{\phi}}\ \sum_{\mathbf{x}}\ \frac{1}{S}\sum_{s=1}^{S}\log q_{\bm{\phi}}\!\big(\mathbf{z}^{(s)}\mid\mathbf{h}\big).$
-2.5 Quantifying Dispersion
-Given a new test prompt $x_{1:T}$, we compute its representation $\mathbf{h}=f_{\theta}(x_{1:T})$ in a single forward pass of the base model.
-The student then predicts the parameters of a $K$-component diagonal-covariance Gaussian mixture over semantic embeddings.
-We measure dispersion using the order-2 Rényi entropy (Rényi, 1961) of the predictive density:
-$\mathsf{H}_{2}\!\big(q_{\phi}(\cdot\mid\mathbf{h})\big)=-\log\int q_{\phi}(\mathbf{z}\mid\mathbf{h})^{2}\,d\mathbf{z},$
-Intuitively, the integral $\int q(\mathbf{z})^{2}d\mathbf{z}$ is the collision probability of the density: it is larger when mass is concentrated, and smaller when the distribution is spread out (Principe, 2010).
+$\max_{\bm{\phi}}\ \sum_{x}\ \frac{1}{S}\sum_{s=1}^{S}\log q_{\bm{\phi}}\!\big(\mathbf{z}^{(s)}\mid\mathbf{h}(x)\big).$
+2.2 Extracting Uncertainty Signals
+At inference time, SSD produces a prompt-conditioned semantic density $q_{\phi}(\mathbf{z}\mid\mathbf{h})$ over answer embeddings, which we treat as a general-purpose uncertainty object.
+In this work, we focus on two primary operations derived from this object: prompt-level dispersion and answer-level likelihood.
+We also highlight that the explicit density enables additional distributional operations such as sampling.
+(i) Prompt evaluation. We use the entropy of the predicted semantic distribution $q_{\phi}(\cdot\mid\mathbf{h})$ to provide a prompt-level uncertainty estimate in a single forward pass.
+(ii) Answer evaluation.
+Given any candidate response $\mathbf{y}^{\prime}$, we embed it as $\mathbf{z}^{\prime}=\Phi(\mathbf{y}^{\prime})$ and score its compatibility with the prompt via the posterior log-likelihood $\log q_{\phi}(\mathbf{z}^{\prime}\mid\mathbf{h})$.
+We may treat answers with low likelihood as out-of-domain relative to the prompt context.
+(iii) Latent sampling.
+Because $q_{\phi}(\mathbf{z}\mid\mathbf{h})$ is an explicit mixture distribution, we can compute summary statistics such as the mixture mean $\mathbb{E}_{q_{\phi}}[\mathbf{z}]=\sum_{k}\pi_{k}(\mathbf{h})\,\bm{\mu}_{k}(\mathbf{h})$ and, when desired, draw latent samples $\tilde{\mathbf{z}}\sim q_{\phi}(\cdot\mid\mathbf{h})$.
+We investigate operations (i) and (ii) in Sections 4.2 and 4.3, respectively.
+Operation (iii) illustrates a structural consequence of modelling the full predictive density and is discussed in Appendix B.2.
+2.2.1 Quantifying Dispersion
+We measure dispersion using the order-2 Rényi entropy [renyi1961measures] of the predictive density $q_{\phi}(\cdot\mid\mathbf{h})$:
+$\mathsf{H}_{2}\!\big(q_{\phi}(\cdot\mid\mathbf{h})\big)=-\log\int q_{\phi}(\mathbf{z}\mid\mathbf{h})^{2}\,d\mathbf{z}.$
+Intuitively, the integral $\int q(\mathbf{z})^{2}d\mathbf{z}$ is the collision probability of the density: it is larger when mass is concentrated, and smaller when the distribution is spread out [principe2010information].
 Higher $\mathsf{H}_{2}$ corresponds to greater semantic dispersion.
-We use Rényi-2 because it yields an exact closed form for Gaussian mixtures (Wang et al., 2009), enabling a sampling-free, analytic dispersion score at inference time.
-By contrast, the Shannon differential entropy (Shannon, 1948) does not, in general, admit a closed-form expression for Gaussian mixture models; it is typically approximated using Monte Carlo estimation or via analytic bounds (Huber et al., 2008; Dahlke and Pacheco, 2023).
+We use Rényi-2 because it yields an exact closed form for Gaussian mixtures [wang2009closed], enabling a sampling-free, analytic dispersion score at inference time.
+By contrast, the Shannon differential entropy [shannon1948mathematical] does not, in general, admit a closed-form expression for Gaussian mixture models; it is typically approximated using Monte Carlo estimation or via analytic bounds [huber2008entropy, dahlke2023closedform].
 Using Rényi-2 therefore preserves SSD’s goal of avoiding additional sampling or expensive numerical estimation at deployment.
-We provide the closed form expression and derivation in Appendix Section A.
-2.6 Using the Predicted Semantic Distribution
-At inference time, SSD produces a prompt-conditioned semantic density $q_{\phi}(\mathbf{z}\mid\mathbf{h})$ over answer embeddings, which we treat as a general-purpose uncertainty object with three primitive operations:
-(i) Semantic Dispersion Prediction
-We use the Rényi-2 entropy $\mathsf{H}_{2}\!\big(q_{\phi}(\cdot\mid\mathbf{h})\big)$ to provide a prompt-level uncertainty estimate in a single forward pass.
-(ii) Context Verification
-Given any candidate response $\mathbf{y}^{\prime}$, we embed it as $\mathbf{z}^{\prime}=\Phi(\mathbf{y}^{\prime})$ and score its compatibility with the prompt via the posterior log-likelihood $\log q_{\phi}(\mathbf{z}^{\prime}\mid\mathbf{h})$.
-We treat answers with low likelihood as out-of-domain relative to the prompt context.
-(iii) Latent Sampling
-Because $q_{\phi}(\mathbf{z}\mid\mathbf{h})$ is an explicit mixture distribution, we can compute summary statistics such as the mixture mean $\mathbb{E}_{q_{\phi}}[\mathbf{z}]=\sum_{k}\pi_{k}(\mathbf{h})\,\bm{\mu}_{k}(\mathbf{h})$ and, when desired, draw latent samples $\tilde{\mathbf{z}}\sim q_{\phi}(\cdot\mid\mathbf{h})$.
+We provide the closed form expression and derivation in Appendix A.
+2.3 Extension to Generic Sequence Domains
+Although we focus on autoregressive language models in this work, the SSD construction is agnostic to modality.
+In language tasks, $\mathbf{z}=\Phi(y_{1:L})$ captures the semantic content of a generated answer.
+More generally, $\mathbf{z}$ may represent any task-relevant latent characterization of a predicted sequence $y_{1:L}$, rather than its individual tokens or events.
+Under stochastic generation, a sequence model induces a predictive distribution over such latent representations.
+SSD distils this sampled distribution into a lightweight conditional density $q_{\phi}(\mathbf{z}\mid\mathbf{h})$, where $\mathbf{h}$ encodes the input sequence.
+The resulting density provides a sampling-free approximation to the model’s belief over high-level outcomes.
+We provide further discussion of potential application domains in Section 5.
+3 Related Work
+Uncertainty estimation in language models.
+xiong2024efficient categorize UQ methods for LLMs into three broad categories: single-sample, multi-sample, and probing-based.
+Single-sample methods are training-free and include output perplexity, negative sequence probability, and mean token entropy [malinin2021uncertainty].
+Probing-based methods similarly use a single sample at inference time but train a predictive model on internal LLM activations to predict an uncertainty-relevant quantity, such as the probability of correctness [kadavath2022languagemodelsmostlyknow] or semantic dispersion [kossen2024semanticentropyprobesrobust].
+Multi-sample methods often rely on notions of semantic dispersion, either explicitly, such as Semantic Entropy [farquhar_detecting_2024] and Number of Clusters [lin2024generating], or implicitly, as with P(true) [kadavath2022languagemodelsmostlyknow].
+We emphasize that these approaches typically model a proxy for language model uncertainty, geared for application in a particular task.
+There is limited consensus on how to decompose classical notions of aleatoric and epistemic uncertainty in LLMs, and the resulting uncertainty signals are often only modestly correlated [hou2024decomposing, xiong2024efficient].
+Distillation in language models.
+Distillation typically refers to training smaller language models on the logit or token outputs of more capable models, with notable success in reasoning domains [guo2025deepseek].
+Such approaches transfer predictive behaviour but do not explicitly target calibrated uncertainty or structured output distributions beyond the token level.
+In contrast, we focus on sequence-level distributions of model outputs and investigate whether these can be elicited without autoregressive sampling.
+piskorz2026eliciting study a related but narrower phenomenon concerning language model predicted distributions over numerical targets, whereas we consider broader semantic targets.
+Knowledge distillation for uncertainty estimation.
+A related line of work seeks to distil ensemble-based uncertainty estimates into a single model [Malinin2020Ensemble, fathullah2022self, landgraf2024dudes].
+Ensembles are widely regarded as producing high-quality uncertainty estimates, but incur substantial computational cost at inference time.
+Recent approaches therefore train a student model to approximate ensemble predictive distributions or uncertainty signals, aiming to recover ensemble-quality uncertainty in a single forward pass, typically by matching logits, predictive variances, or other distributional summaries [fathullah2023logit, park2025knowledge].
+In contrast, our setting does not assume access to an external ensemble and does not aim to replicate variance estimates or token-level predictive distributions.
+Instead, we distil a language model’s semantic sequence-level output distribution induced by stochastic sampling, learning a prompt-conditioned density over answer embeddings.
+The most closely related approaches in the LLM literature are probing-based uncertainty estimators [kadavath2022languagemodelsmostlyknow, kossen2024semanticentropyprobesrobust], which however compress uncertainty into a single scalar statistic.
+By contrast, we predict a full sequence-level distribution in a single forward pass, enabling downstream operations beyond scalar risk scoring, including likelihood-based post-generation verification and multiple-choice answer selection.
+Rather than estimating a task-specific uncertainty heuristic, we approximate the underlying semantic output distribution itself.
+Figure 2: Qualitative analysis of the SSD density.
+(A) A correct answer yields minimal SSD entropy (0.1st percentile), while a hallucination yields high uncertainty (99.7th), aligning with sampling-based Teacher Dispersion (TD) and Semantic Entropy (SE).
+Percentiles denote rank within the TriviaQA test set (N=1, [^0] N{=}1{,}000).
+(B) A precise prompt (P1) yields significantly lower entropy than an underspecified one (P2).
+(C) The SSD likelihood enables context-aware scoring: given identical candidate sets, the student assigns the highest likelihood to the semantically appropriate answer for each prompt.
 4 Experiments
+We evaluate SSD across multiple LLM families in order to answer two core questions.
+First, does SSD faithfully distil sampling-based semantic dispersion into a single-pass model?
+Second, what additional capabilities arise from modelling a full sequence-level density rather than a scalar uncertainty proxy?
+To address the first question, we study prompt-level hallucination prediction, treating the student’s entropy as a pre-generation risk score and comparing it against sampling-based and probing baselines.
+To address the second, we evaluate answer-level likelihood scores, demonstrating how the predicted density supports post-generation verification, including out-of-domain answer detection and multiple-choice answer selection.
+Figure 2 illustrates examples of each operation type.
 4.1 Models and Dataset Generation
-We evaluate the SSD distribution in the context of short-form question answering, using the TriviaQA dataset (Joshi et al., 2017).
-We construct a training set of $N_{train}=4000$ prompts and report results on a held-out test set of $N_{test}=1000$ prompts.
-We evaluate across five model families: Mistral (Ministral-8B-Instruct), Llama-3 (3.1-8B-Instruct, 3.2-3B-Instruct), Qwen (Qwen3-4B-Instruct-2507, Qwen3-8B), Gemma (Gemma-3-4B) and SmolLM (SmolLM3-3B).
-For all experiments, we use Qwen3-8B as the oracle ‘judge’ model to determine the correctness of generated answers.
-For each prompt and model, we sample $S=32$ samples at temperature $T=1$.
-We additionally sample a single ‘default’ answer at $T=0.1$ which we input to the judge model to generate the hallucination label.
-For the answer embedding, we use EmbeddingGemma (Vera et al., 2025), with the embedding truncated to a dimension of $d_{z}=128$ .
-For hallucination detection, we further reduce dimensionality using PCA to $d_{z}^{\prime}=16$, following ablation studies detailed in Appendix Section D.
+We evaluate SSD in the context of short-form question answering, using the TriviaQA [joshi-etal-2017-triviaqa] and MMLU [hendrycks2021measuring] datasets.
+We construct training sets of $N_{train}=4{,}000$ prompts and report results on held-out test sets of $N_{test}=1{,}000$ prompts.
+MMLU is formatted as a multiple-choice benchmark, as opposed to the open-ended TriviaQA.
+To properly assess semantic dispersion and enable comparison across datasets, we use a filtered subset of MMLU containing only open-ended questions [myrzakhan2024openllmleaderboardmultichoiceopenstylequestions].
+We select these datasets because they represent two distinct task types: knowledge-seeking (TriviaQA) and reasoning-heavy (MMLU).
+Prior work has found semantic dispersion to be a useful proxy for language model uncertainty in the former task type and less so in the latter [xiong2024efficient]; as such MMLU provides a more challenging setting for evaluating SSD, although in practice it reflects a mixture of both task types.
+We evaluate across five model families: Mistral (Ministral-8B-Instruct), Llama-3 (3.1-8B-Instruct, 3.2-3B-Instruct), Qwen (Qwen3-4B-Instruct-2507, Qwen3-8B), Gemma (Gemma-3-4B) and SmolLM (SmolLM3-3B) [grattafiori2024llama, mistralai_ministraux_2024, yang2025qwen3technicalreport, gemmateam2025gemma3technicalreport, bakouch2025smollm3].
+These models span multiple architectural families, parameter scales (3B-8B), and training pipelines, allowing us to assess the robustness and generality of SSD across diverse backbone designs rather than tailoring it to a single model.
+For all experiments, we use Qwen3-8B as the oracle ‘judge’ model to determine the correctness of generated answers given a reference answer [zheng2023judge].
+For each prompt and model, we generate $S=32$ stochastic samples at temperature $T=1$.
+We additionally sample a single ‘default’ answer at $T=0.1$, which is provided to the judge model to obtain the hallucination label.
+For answer embeddings, we use EmbeddingGemma [vera2025embeddinggemmapowerfullightweighttext], with the embedding truncated to a dimension of $d_{z}=128$.
+For hallucination prediction, we further reduce dimensionality using PCA to $d_{z}^{\prime}=16$, following ablation studies detailed in Appendix C.
 For likelihood-based OOD verification, we set $d_{z}^{\prime}=32$.
-3.2 Layer Selection Protocol
-To estimate a suitable layer $\ell^{*}$ for the student input, we use the results of our hallucination detection baseline experiments, described in Section 3.3.
-Following Kossen et al.
-(2024), we first train simple linear probes on all layers to predict the binarized Semantic Entropy.
-We select the layer index $\ell^{*}$ that maximizes validation accuracy for this proxy task.
-We assume that the layer most predictive of the scalar entropy value also contains the richest information for constructing the full distributional mixture.
-We find this typically corresponds to deeper layers, shown in Table 5 in the Appendix.
-3.3 Hallucination Detection via Predicted Dispersion
-We treat the student’s dispersion score as a pre-generation risk score: higher values indicate greater semantic spread and higher hallucination risk.
+4.2 Prompt-Level Hallucination Prediction
+Table 1: Hallucination prediction trade-offs.
+AUROC is reported per dataset as a macro-average (mean $\pm$ std) across models.
+Latency summarizes inference-time requirements (single-pass vs.
+$S$ sampled generations, with entailment-based SE additionally requiring $S^{2}$ natural language inference (NLI) comparisons).
+Target cost denotes the computation required to construct supervision targets during training; $C_{CL}$ is the cost of acquiring correctness labels.
+Only SSD produces an explicit predictive density, enabling post-generation likelihood scoring and latent sampling in addition to pre-generation uncertainty.
+Method TriviaQA $\uparrow$ MMLU $\uparrow$ Latency $\downarrow$ Target cost $\downarrow$ Predictive density
+SE 0.804 $\pm$ 0.035 0.663 $\pm$ 0.040 $S$ samples + $S^{2}$ NLI $\times$
+SEP 0.759 $\pm$ 0.026 0.628 $\pm$ 0.029 Single-pass $S$ samples + $S^{2}$ NLI $\times$
+PCP 0.771 $\pm$ 0.020 0.679 $\pm$ 0.019 Single-pass $C_{\text{CL}}$ $\times$
+TD 0.699 $\pm$ 0.027 0.647 $\pm$ 0.023 $S$ samples $\times$
+SSD 0.708 $\pm$ 0.024 0.623 $\pm$ 0.025 Single-pass $S$ samples + $S\,\Phi$ $\checkmark$
+Table 2: Hallucination prediction, comparison with teacher dispersion.
+TD computes dispersion from $S=32$ sampled semantic embeddings at inference time, while SSD predicts a prompt-conditioned semantic density and scores dispersion analytically.
+We report hallucination prediction AUROC/AUPRC in % as the mean over 1,000 bootstrap resamples of the test set, with standard deviation as subscript; bold indicates the better of TD vs.
+SSD per metric and dataset.
+AUROC AUPRC
+Model TD SSD TD SSD
+TriviaQA
+Qwen3 8B 68.31.6 67.71.8 52.42.6 56.92.8
+Qwen3 4B 73.21.6 69.91.6 59.72.4 61.92.5
+Llama 3.1 8B 69.61.7 68.71.9 44.42.7 51.92.9
+Llama 3.2 3B 66.61.7 73.71.6 55.82.6 65.52.5
+Ministral 8B 74.31.7 74.91.6 57.92.8 59.92.8
+SmolLM3 3B 69.91.6 71.11.6 57.12.5 62.42.5
+Gemma 3 4B 67.61.6 70.01.7 63.12.3 66.12.5
+MMLU
+Qwen3 8B 63.51.7 57.11.8 61.92.3 55.92.3
+Qwen3 4B 63.11.8 64.91.7 58.52.4 59.42.5
+Llama 3.1 8B 69.01.6 63.81.7 70.92.1 62.72.4
+Llama 3.2 3B 67.11.8 63.81.7 74.02.0 74.51.9
+Ministral 8B 65.11.8 63.61.8 71.42.1 71.72.1
+SmolLM3 3B 62.31.8 62.01.8 64.62.2 63.02.3
+Gemma 3 4B 62.81.7 60.71.8 71.12.0 69.22.1
+We first evaluate whether the distilled semantic density provides a reliable pre-generation hallucination signal.
+Specifically, we treat the entropy of the predicted distribution as a prompt-level risk score: higher entropy corresponds to greater semantic dispersion and increased hallucination risk.
 We report AUROC and AUPRC against correctness labels determined using an LLM-as-judge framework.
 SSD compares favourably to teacher dispersion.
-We first compare the student’s dispersion score to the standard deviation of the teacher’s $S=32$ semantic embeddings, which we term teacher dispersion (TD).
-Table 1 reports results across models.
-Despite requiring no inference-time sampling, SSD outperforms TD on four of the seven models investigated in AUROC and on all seven in AUPRC.
+We compare the student’s dispersion score to the standard deviation of the teacher’s $S=32$ semantic embeddings, which we term teacher dispersion (TD).
+Table 2 reports results across models and datasets.
+On the TriviaQA dataset, despite requiring no inference-time sampling, SSD outperforms TD on four of the seven models investigated in terms of AUROC and on all seven in AUPRC.
 We attribute this in part to a smoothing effect: TD is a finite-sample statistic and assigns zero uncertainty to prompts where the model consistently produces the same answer, even when that answer is incorrect.
 By contrast, SSD learns a continuous mapping from prompt representations to semantic uncertainty, enabling differentiated risk estimates in such degenerate cases, which likely contributes to its stronger AUPRC.
-Model AUROC AUPRC
-TD SSD TD SSD
-Qwen3-8B 0.684 $\pm$ 0.017 0.677 $\pm$ 0.018 0.525 $\pm$ 0.027 0.570 $\pm$ 0.028
-Qwen3-4B 0.732 $\pm$ 0.016 0.699 $\pm$ 0.016 0.596 $\pm$ 0.025 0.618 $\pm$ 0.025
-Llama 3.1 8B 0.695 $\pm$ 0.017 0.686 $\pm$ 0.020 0.443 $\pm$ 0.028 0.519 $\pm$ 0.030
-Llama 3.2 3B 0.665 $\pm$ 0.017 0.738 $\pm$ 0.017 0.558 $\pm$ 0.027 0.657 $\pm$ 0.027
-Ministral 8B 0.743 $\pm$ 0.016 0.748 $\pm$ 0.016 0.579 $\pm$ 0.029 0.599 $\pm$ 0.029
-SmolLM3 3B 0.698 $\pm$ 0.016 0.711 $\pm$ 0.016 0.569 $\pm$ 0.025 0.623 $\pm$ 0.026
-Gemma 3 4B 0.676 $\pm$ 0.017 0.701 $\pm$ 0.017 0.632 $\pm$ 0.024 0.662 $\pm$ 0.025
-Table 1: Distilling sampled dispersion into a single pass.
-TD computes dispersion from $S=32$ sampled semantic embeddings at inference time, while SSD predicts a prompt-conditioned semantic density and scores dispersion analytically.
-We report hallucination detection AUROC/AUPRC as mean $\pm$ standard deviation over 1,000 bootstrap resamples of the test set; bold indicates the better of TD vs.
-SSD per metric.
+On the MMLU dataset, SSD is a relatively weaker performer compared to the sampled teacher signal, which we attribute to the task being less suited for the semantic dispersion uncertainty proxy.
+SSD still surpasses TD in three out of seven models from an AUPRC perspective.
+Overall, these results demonstrate that SSD successfully distils sampling-based semantic dispersion into a single-pass density estimator.
+Despite requiring no inference-time sampling, the student recovers much of the teacher’s uncertainty signal, and in some cases improves upon the finite-sample baseline.
 Comparison to additional baselines.
-We further compare SSD against three widely used hallucination detection methods:
-  1. 1.
-    Probability of Correctness Probe (PCP): A logistic regression classifier trained on the prompt representation to predict binary correctness.
-Originally proposed as ‘P(I know)’ by Kadavath et al.
-(2022), this method requires externally provided correctness labels.
-  2. 2.
-    Semantic Entropy (SE): A state-of-the-art sampling-based approach that measures semantic dispersion by clustering $S$ stochastic completions using a natural language inference (NLI) model to assess mutual entailment (Farquhar et al., 2024).
-  3. 3.
-    Semantic Entropy Probe (SEP): A lightweight probe with the same architecture as PCP, trained to regress a binarized SE target, inheriting SE’s sampling and NLI-based target construction cost (Kossen et al., 2024).
-To reflect agentic deployment constraints, we report hallucination detection performance alongside method characteristics: inference latency, supervision target construction cost, and whether a method supports post-generation (posterior) signals in addition to pre-generation (prior) uncertainty.
-Table 2 summarizes these trade-offs and reports macro-average AUROC across model families; full per-model AUROC/AUPRC tables are provided in the Appendix (tables 8 and 9).
-Overall, SE achieves the strongest AUROC, consistent with prior work showing that explicit sampling of multiple reasoning paths yields high-quality uncertainty signals (Kadavath et al., 2022).
-Among single-pass methods, PCP and SEP outperform SSD as pre-generation hallucination predictors, confirming that the binary classification problem is simpler than distilling the geometry of a full semantic distribution.
-However, SSD occupies a distinct and practically useful regime: it matches probe-level inference latency, avoids entailment-based target construction, and uniquely provides posterior signals as well as prior risk scores.
-The post-generation signals enable verification and filtering mechanisms that are not available to scalar probes.
-Method AUROC $\uparrow$ Latency $\downarrow$ Target cost $\downarrow$ Posterior
-SE 0.804 $\pm$ 0.036 $S$ samples + $S^{2}$ NLI $\times$
-SEP 0.759 $\pm$ 0.026 Single-pass $S$ samples + $S^{2}$ NLI $\times$
-PCP 0.771 $\pm$ 0.021 Single-pass $C_{\text{CL}}$ $\times$
-TD 0.699 $\pm$ 0.027 $S$ samples $\times$
-SSD 0.708 $\pm$ 0.024 Single-pass $S$ samples + $S\,\Phi$ $\ell(y),\ \mathbb{E}[\mathbf{z}]$
-Table 2: Hallucination detection trade-offs.
-AUROC is reported as a macro-average (mean $\pm$ std) across models.
-Latency summarizes inference-time requirements (single-pass vs $S$ sampled generations, with entailment-based SE additionally requiring $S^{2}$ natural language inference (NLI) comparisons).
-Target cost denotes the computation required to construct supervision targets during training; $C_{CL}$ is the cost of acquiring correctness labels.
-Only SSD explicitly provides post-generation posterior signals in addition to pre-generation uncertainty.
-3.4 Distillation Fidelity vs. Detection Performance
-We examine whether the quality of distillation (how closely the student learns the teacher distribution) constrains hallucination detection performance.
-We define distillation fidelity ($\rho_{\text{fidelity}}$) as the Spearman rank correlation, over $N=1000$ test prompts, between the student’s predicted entropy and the teacher dispersion.
-A high $\rho_{\text{fidelity}}$ indicates a more faithful recovery of the teacher’s prompt-induced uncertainty ordering.
-Figure 2 plots hallucination detection AUROC against distillation fidelity across base models and student hyperparameter settings.
-We observe a strong positive relationship, with a cross-model Spearman correlation of $\rho_{\text{meta}}=0.65$, indicating that detection performance is limited by the student’s ability to learn the prompt-to-dispersion mapping.
-The failure cases, notably Qwen3-4B, exhibit near-zero $\rho_{\text{fidelity}}$, suggesting that the relevant uncertainty signal is not accessible from the chosen prompt representation.
-Figure 2: Distillation fidelity drives detection performance.
-The X-axis shows distillation fidelity ($\rho_{\text{fidelity}}$): the Spearman correlation between the student’s predicted entropy and the teacher dispersion across test prompts.
-Multiple points per model correspond to different student hyperparameter settings.
-$\rho_{\text{meta}}$ is the Spearman correlation between the X and Y variables; it shows that models where the student fails to learn the distribution (low $\rho_{\text{fidelity}}$) result in poor detection.
-3.5 Context Verification via Out-of-Domain Detection
-We evaluate the SSD density as a post-generation verification signal by measuring its ability to detect mismatched prompt-answer pairs.
+We compare SSD against three widely used hallucination prediction methods: (1) Probability of Correctness Probe (PCP), a logistic regression classifier trained on the prompt representation to predict correctness.
+Originally proposed as ‘P(I know)’ by kadavath2022languagemodelsmostlyknow, this method requires externally provided correctness labels.
+(2) Semantic Entropy (SE), a sampling-based approach that measures semantic dispersion by clustering $S$ completions using a natural language inference (NLI) model to assess mutual entailment [farquhar_detecting_2024].
+(3) Semantic Entropy Probe (SEP), a lightweight probe with the same architecture as PCP, trained to regress a binarized SE target, and inheriting SE’s sampling and NLI-based target construction cost [kossen2024semanticentropyprobesrobust].
+To provide a fuller comparison of methods, we report hallucination prediction performance alongside method characteristics: inference latency, training supervision target construction cost, and whether a method supports post-generation signals in addition to pre-generation uncertainty.
+Table 1 summarizes these trade-offs and reports macro-average AUROC across model families; full per-model AUROC/AUPRC tables are provided in Appendix F.
+Overall, SE achieves the highest AUROC on TriviaQA, whilst PCP is the strongest performer on MMLU, consistent with prior work [xiong2024efficient].
+While SE and PCP achieve the best performance in their respective regimes, these methods either incur significant sampling cost (SE) or reduce uncertainty to a supervised scalar classification problem (PCP).
+SSD occupies a distinct regime: it matches probe-level inference latency, avoids entailment-based target construction during training, and preserves an explicit predictive density.
+Although its pre-generation hallucination prediction performance is slightly weaker than specialized probes, it uniquely provides post-generation likelihood signals derived from the same model.
+We further observe that the SSD performance is model-dependent.
+On TriviaQA, for example, it achieves an AUROC of 74.9 for Ministral 8B but only 67.7 for Qwen3 8B (Table A5).
+Analysis in Appendix B.1 indicates that this variation is driven primarily by the fidelity of the distillation process: for some models, it is particularly difficult to predict the semantic distribution from the prompt representation.
+We discuss these findings further in Section 5.
+4.3 Answer-Level Response Evaluation
+Having established that SSD recovers sampling-based dispersion signals, we next evaluate the additional capabilities enabled by modelling an explicit predictive density.
+Unlike scalar uncertainty scores, the density allows us to assess the quality of specific candidate answers by measuring their likelihood under the model’s predicted semantic distribution.
+We examine this utility in two settings: out-of-domain answer detection (binary classification) and multiple-choice question answering (multi-class selection).
+4.3.1 Out-of-domain Answer Detection
+We first evaluate the student’s ability to detect mismatched prompt-answer pairs.
 We frame this as an out-of-domain detection task: for each test prompt $x_{i}$, we pair its representation $\mathbf{h}_{i}$ with (i) the model’s default answer $y_{i}$ and (ii) a random answer $y_{j}$ drawn from a different prompt.
-Each candidate answer is embedded as $\mathbf{z}=\Phi(y)$ and scored using the student’s posterior log-likelihood $\log q_{\phi}(\mathbf{z}\mid\mathbf{h})$.
+Each candidate answer is embedded as $\mathbf{z}=\Phi(y)$ and scored using the student’s log-likelihood $\log q_{\phi}(\mathbf{z}\mid\mathbf{h})$.
 We report AUROC for distinguishing matched from mismatched pairs.
-As shown in Table 3, the posterior likelihood is a strong context-verification signal, achieving AUROCs above 0.95 for most base models.
-This suggests that SSD can serve as a low-latency guardrail in agentic workflows, for example detecting cache poisoning (Chen et al., 2024) or filtering retrieved content in retrieval-augmented generation (Lewis et al., 2020).
-The likelihood provides a modest signal for post-generation hallucination detection, demonstrated by AUROCs above 0.5.
-This result could be interpreted through the lens of Semantic Density (Qiu and Miikkulainen, 2024), which posits that reliable samples tend to lie in denser regions of a semantic predictive space.
-Rather than approximating this distribution via samples and associated sequence likelihoods, our approach explicitly models it.
-Table 3: Utility of SSD posterior likelihood.
-Out-of-domain (OOD) detection reports AUROC for distinguishing an answer belonging to the prompt from a random answer.
-Hallucination detection compares AUROC for pre-generation entropy vs.
-post-generation likelihood.
-While likelihood is a strong OOD detector, entropy remains the superior signal for hallucination classification.
+As shown in Table 3, the SSD log-likelihood is a strong context-verification signal, achieving AUROCs above 0.95 for most base models on both TriviaQA and MMLU.
+Practically, this could enable SSD to act as a content ‘guard’ [inan2023llamaguardllmbasedinputoutput], for example detecting cache poisoning [chenagentpoison2024] or filtering retrieved content in retrieval-augmented generation [lewisrag2020].
+We also evaluate the student likelihood as a post-generation hallucination signal, motivated by prior work showing that probability density in semantic space is strongly predictive of correctness [qiu2024semanticdensity].
+While informative, our likelihood signal is weaker than the pre-generation entropy, with AUROC scores of approximately 0.6.
+Unlike sampling-based density methods, which construct the semantic distribution from multiple generations at inference time, SSD predicts the density from the prompt representation alone.
+The weaker performance reflects the limits of this single-pass approximation to a sampling-derived density.
+Table 3: Out-of-domain answer detection.
+We report AUROC for two binary classification tasks using the student’s log-likelihood score.
+Mismatch measures the ability to distinguish the model’s answer to the prompt from a random answer from the dataset.
+Hallucination measures the ability to distinguish correct model-generated answers from incorrect ones.
 Values report mean AUROC with bootstrap standard deviation over 1,000 resamples.
-Model OOD Hallucination
-Likelihood Entropy Likelihood
-Qwen3 8B 0.983 $\pm$ 0.002 0.697 $\pm$ 0.017 0.571 $\pm$ 0.019
-Qwen3 4B 0.984 $\pm$ 0.002 0.703 $\pm$ 0.016 0.555 $\pm$ 0.018
-Llama 3.1 8B 0.989 $\pm$ 0.002 0.702 $\pm$ 0.019 0.578 $\pm$ 0.020
-Llama 3.2 3B 0.980 $\pm$ 0.003 0.728 $\pm$ 0.016 0.612 $\pm$ 0.018
-Ministral 8B 0.944 $\pm$ 0.005 0.758 $\pm$ 0.016 0.616 $\pm$ 0.018
-SmolLM3 3B 0.988 $\pm$ 0.002 0.676 $\pm$ 0.018 0.563 $\pm$ 0.018
-Gemma 3 4B 0.895 $\pm$ 0.007 0.679 $\pm$ 0.017 0.577 $\pm$ 0.018
-3.6 Semantic Consensus from the Predicted Mixture
-We evaluate the generative utility of the student by measuring its ability to approximate the semantic consensus of the model.
-We treat the empirical mean of $S=32$ stochastic samples as a reference semantic centroid, then measure the mean squared distance (MSD) in Euclidean space between this centroid and both (i) the model’s low-temperature output and (ii) the analytical mean of the student mixture, $\bm{\mu}_{ssd}=\sum_{k}\pi_{k}\bm{\mu}_{k}$.
-Lower MSD indicates better recovery of the model’s aggregate semantic belief.
-MSD is computed as $\mathbb{E}_{i}\left[\|\hat{\mathbf{z}}_{i}-\bm{\mu}^{\text{true}}_{i}\|_{2}^{2}\right]$ over examples $i$ in the subset.
-We report results for the best-performing student configuration (Ministral 8B), reflecting the strong architectural dependence of distillation fidelity.
-As shown in Table 4, the SSD mixture mean provides a closer approximation to the semantic centroid than a single default decoding pass.
-This effect is most pronounced for incorrect answers, where SSD improves centroid recovery in over 76% of test cases.
-This aligns with prior work showing that hallucinated outputs tend to deviate from semantic consensus (Lin et al., 2024; Phillips et al., 2025).
-Together with the OOD results in Table 3, this result suggests that SSD’s pre-generation uncertainty estimates can be complemented by post-generation distance- and likelihood-based checks derived from the same predicted density.
-Table 4: Semantic consensus approximation with Ministral 8B.
-We report the mean squared distance (MSD) of the default answer embedding and the SSD predicted mean to the true semantic centroid (average of $S=32$ samples).
-Values denote mean MSD, with the bootstrap standard deviation expressed as a percentage of the mean shown as a subscript.
-Subset Default MSD $\downarrow$ SSD MSD $\downarrow$ Imp. (%) $\uparrow$ SSD Win Rate (%) $\uparrow$
-All 0.0153.2 0.0081.9 +45.5 58.3
-Correct 0.0134.4 0.0092.3 +32.3 48.8
-Incorrect 0.0184.2 0.0073.4 +63.0 76.5
-4 Future Work
+Model Mismatch Hallucination
+TriviaQA
+Qwen3 8B 0.984 $\pm$ 0.002 0.571 $\pm$ 0.019
+Qwen3 4B 0.984 $\pm$ 0.002 0.557 $\pm$ 0.018
+Llama 3.1 8B 0.989 $\pm$ 0.002 0.579 $\pm$ 0.019
+Llama 3.2 3B 0.980 $\pm$ 0.003 0.613 $\pm$ 0.017
+Ministral 8B 0.945 $\pm$ 0.005 0.616 $\pm$ 0.017
+SmolLM3 3B 0.988 $\pm$ 0.002 0.562 $\pm$ 0.018
+Gemma 3 4B 0.895 $\pm$ 0.007 0.578 $\pm$ 0.018
+MMLU
+Qwen3 8B 0.975 $\pm$ 0.003 0.526 $\pm$ 0.018
+Qwen3 4B 0.960 $\pm$ 0.004 0.566 $\pm$ 0.019
+Llama 3.1 8B 0.981 $\pm$ 0.003 0.576 $\pm$ 0.018
+Llama 3.2 3B 0.942 $\pm$ 0.005 0.519 $\pm$ 0.019
+Ministral 8B 0.966 $\pm$ 0.004 0.529 $\pm$ 0.019
+SmolLM3 3B 0.985 $\pm$ 0.003 0.528 $\pm$ 0.018
+Gemma 3 4B 0.923 $\pm$ 0.006 0.548 $\pm$ 0.019
+4.3.2 Multiple-choice Answer Selection
+We next evaluate whether the student’s predicted density can be used to identify the correct answer in a multiple-choice setting.
+We utilize the open-ended subset of MMLU provided by myrzakhan2024openllmleaderboardmultichoiceopenstylequestions, which filters for questions that are self-contained and do not require the context of multiple-choice options to answer.
+The student is trained using only open-ended completions to prompts in the training set.
+For the test set, we take the four provided options ($A,B,C,D$) for each question, embed them using $\Phi$, and rank them according to their log-likelihood $\log q_{\phi}(\mathbf{z}^{\prime}\mid\mathbf{h})$ under the student mixture.
+Table 4 reports the accuracy of selecting the highest-likelihood option (Top-1) alongside the base model’s open-ended answer accuracy as assessed by an LLM-as-judge.
+SSD consistently exceeds the 0.25 random baseline, and for our best performing models (Llama 3.2 and Ministral) approaches the base model’s open-ended accuracy.
+The open-ended accuracy of the base model provides a rough estimate of its dataset-specific knowledge.
+Despite being trained only on sampled open-ended completions and without multiple-choice supervision, the student approaches this level of performance when provided with the options, suggesting that the distilled semantic density captures meaningful structure in the model’s answer distribution.
+Table 4: Multiple-choice answer selection.
+We evaluate the ability of the SSD student - trained solely on open-ended generations - to identify the correct answer in a multiple-choice setting using an open-ended subset of MMLU.
+We report Top-1 accuracy, where the option with the highest predicted likelihood is selected (random baseline =0.25=0.25).
+Open-Ended accuracy denotes the base model’s accuracy when generating answers directly, as determined by an LLM-as-judge.
+Model SSD Top-1 Open-Ended
+Qwen3 8B 0.327 $\pm$ 0.015 0.515 $\pm$ 0.015
+Qwen3 4B 0.312 $\pm$ 0.014 0.545 $\pm$ 0.016
+Llama 3.1 8B 0.339 $\pm$ 0.015 0.483 $\pm$ 0.015
+Llama 3.2 3B 0.346 $\pm$ 0.015 0.358 $\pm$ 0.015
+Ministral 8B 0.363 $\pm$ 0.015 0.396 $\pm$ 0.015
+SmolLM3 3B 0.332 $\pm$ 0.015 0.466 $\pm$ 0.016
+Gemma 3 4B 0.292 $\pm$ 0.014 0.399 $\pm$ 0.015
+Summary of experimental findings.
+Taken together, the prompt-level and answer-level evaluations demonstrate that SSD’s value lies not in a particular scalar uncertainty score, but in exposing an explicit semantic belief state from which multiple reliability signals can be derived.
+By modelling a prompt-conditioned distribution rather than a scalar proxy, SSD unifies pre-generation risk estimation and post-generation verification within a single, low-latency framework.
+The performance variation observed across base LLMs points to differences in how semantic uncertainty is encoded in internal activations.
+5 Discussion & Future Work
+The experimental results highlight both the strengths and current limitations of SSD.
+While modelling a full semantic density enables richer post-generation capabilities than scalar probes, its effectiveness depends on how faithfully the student can reconstruct the teacher’s sampled distribution from prompt representations alone.
+We now discuss these limitations and outline directions for strengthening the distributional framework.
 Improving distillation fidelity.
-Our results indicate that SSD’s utility as a pre-generation hallucination signal is primarily limited by distillation fidelity: detection performance improves when the student accurately recovers the teacher’s prompt-induced uncertainty ordering.
-Promising directions include scaling the distillation dataset, enriching the student input beyond a single-token/single-layer representation, and distilling from stronger teachers (larger $S$, or higher-quality semantic targets).
-These extensions would clarify whether observed failures reflect an information bottleneck in the prompt representation or insufficient supervision for density estimation.
-Combining prior and posterior signals.
-SSD exposes complementary reliability signals from a single predicted density: pre-generation dispersion forecasts risk, while post-generation likelihood and distance-to-consensus verify prompt–answer agreement.
-A natural next step is to integrate these signals into decision policies for agents, for example abstaining or retrieving evidence under high dispersion, and accepting or rejecting candidate actions based on likelihood or consensus distance.
+Our results in Appendix B.1 show a strong correlation between distillation fidelity and hallucination prediction performance, indicating that SSD’s effectiveness depends on how accurately the student reconstructs the teacher’s prompt-induced uncertainty ordering.
+Promising directions to improve this fidelity include scaling the distillation dataset, enriching the student input beyond a single-token/single-layer representation, and distilling from stronger teachers (for example sampling more responses per question, or using higher-quality semantic targets).
+Subsequent findings may clarify whether observed performance gaps between models stem from limited semantic information in the prompt representation or insufficient supervision for learning high-dimensional densities.
+Combining pre- and post-generation signals.
+SSD exposes complementary reliability signals from a single predicted density: pre-generation dispersion forecasts risk, while post-generation likelihood verifies prompt–answer compatibility.
+A natural next step is to integrate these signals into unified decision policies for LLM-based systems, for example abstaining or retrieving evidence under high dispersion, and accepting or rejecting candidate actions based on likelihood thresholds.
 Diffusion models.
-While we focus on autoregressive models, SSD may be particularly well suited to diffusion-based language models, where hidden prompt representations may more directly parameterize the final semantic output distribution (Sahoo et al., 2024).
+Although our experiments focus on autoregressive models, SSD may be particularly well suited to diffusion-based language models [sahoodiffusion2024].
+Since diffusion models do not rely on token-wise autoregressive decoding, the mapping from prompt representation to final output distribution may be more direct, potentially reducing the gap between the model’s internal belief state and the semantic distribution induced by sampling.
 Domain-specific uncertainty estimation.
-The technique naturally extends to domain-specific foundation models, such as those in healthcare (Renc et al., 2024; Steinberg et al., 2024) and time-series forecasting (Faw et al., 2025).
-By distilling predictive distributions over latent outcome representations, students could provide real-time uncertainty estimates for future patient trajectories or other temporal processes, enabling risk-aware decision-making in sequential settings (Zeng et al., 2025).
+As described in Section 2.3, SSD extends beyond language to sequence models whose outputs admit meaningful latent representations.
+We see particular potential utility for this technique in electronic health record (EHR) foundation models, which tokenize longitudinal patient histories and generate future clinical event sequences [renc2024zero].
+These models typically predict discrete events, but clinical utility often depends on a higher-level characterization of the patient’s future state, such as overall disease burden or risk profile [steinberg2024motor, zeng2025trajsurvlearningcontinuouslatent].
+In this setting, the embedding model could correspond to a transformer-based EHR encoder such as BEHRT [li2020behrt], summarizing a predicted trajectory into a clinically informative latent health representation.
+Sampling from $q_{\phi}(\mathbf{z}\mid\mathbf{h})$ could then yield low-latency forecasts of plausible future patient states, while the entropy of the density would quantify uncertainty over long-term outcomes.
+Similar structure arises in time-series foundation models [faw2025incontext], where future trajectories may be summarized by latent representations capturing regime, trend, or seasonal structure [wang2022learning].
 6 Conclusion
-We introduced Semantic Self-Distillation (SSD), a distributional probing framework that distils a language model’s sampled semantic answer distribution into a lightweight, prompt-conditioned density estimator.
-Unlike scalar probes which compress uncertainty into a single score, SSD predicts an explicit semantic density, enabling analytic dispersion scoring and additional distribution-level operations in a single forward pass.
-Empirically, SSD provides competitive hallucination-risk signals relative to finite-sample baselines on several model families, while avoiding the inference-time overhead of generating multiple completions.
-Crucially, modelling the full density unlocks capabilities that scalar probes cannot match: the posterior likelihood provides strong context verification checks for out-of-domain answers, while mixture statistics allow for single-pass semantic consensus estimates.
-Overall, SSD demonstrates that distilling a semantic predictive distribution can yield a general-purpose uncertainty object with applications in forecasting and verification.
-By decoupling these capabilities from expensive inference-time sampling, SSD offers a practical framework for integrating principled uncertainty quantification into latency-critical agentic applications.
-6 Acknowledgments
+We introduced Semantic Self-Distillation (SSD), a framework for distilling a language model’s sampled semantic answer distribution into a lightweight, prompt-conditioned predictive density.
+Rather than reducing uncertainty to a scalar score, SSD models the geometry of the semantic output space itself, enabling analytic dispersion estimates and likelihood-based evaluation in a single forward pass.
+Empirically, SSD provides competitive hallucination prediction performance relative to sampling-based semantic dispersion baselines across several model families, while avoiding the inference-time overhead of generating multiple completions.
+Crucially, modelling the full density enables capabilities unavailable to scalar probes, including answer-level reliability scoring and distribution-based selection.
+More broadly, SSD demonstrates that uncertainty in complex generative models can be distilled as an explicit density over latent outcome representations.
+By shifting computational cost from inference to training, SSD provides a practical mechanism for integrating principled, distribution-level uncertainty into latency-critical language systems and other sequence domains.
+Acknowledgements.
 DAC was funded by an NIHR Research Professorship; a Royal Academy of Engineering Research Chair; and the InnoHK Hong Kong Centre for Cerebro-cardiovascular Engineering (COCHE); and was supported by the National Institute for Health Research (NIHR) Oxford Biomedical Research Centre (BRC) and the Pandemic Sciences Institute at the University of Oxford.
@@
 This computation is $\mathcal{O}(K^{2})$ in the number of mixture components and is negligible relative to the transformer backbone.
-Appendix B Model Metadata
-Table 5: Model characteristics.
-‘Base Acc’ denotes QA accuracy.
-‘PCP’ and ‘SEP’ indicate optimal probe layers.
-The ‘Best SSD Config’ columns show the student architecture (HH=Hidden, DD=Depth, KK=Components) that achieved the highest SSD AUROC.
-Probe Layers Best SSD Config
-Model Base Acc PCP SEP H D K AUROC
-Qwen3-8B 0.620 32 24 128 6 10 0.677
-Qwen3-4B-Instruct-2507 0.565 18 20 256 2 10 0.699
-Llama 3.1 8B 0.700 20 16 256 4 2 0.687
-Llama 3.2 3B 0.585 15 19 512 3 5 0.737
-Ministral 8B 0.656 33 34 1024 6 1 0.748
-SmolLM3 3B 0.567 20 19 512 6 5 0.710
-Gemma 3 4B 0.532 22 19 512 2 10 0.700
-Appendix C Computational Complexity Analysis
-The primary motivation for Semantic Self-Distillation is to enable generic uncertainty estimation in AR models without sampling.
-Here we compare the asymptotic complexity of our method against sampling-based baselines.
-Let $T_{\text{pre}}$ be the prompt length, $T_{\text{gen}}$ be the generation length, $S$ be the number of samples, and $C_{\text{LLM}}$ be the cost of a single transformer forward pass per token.
-Sampling-based Methods (SE, TD).
-Computing Semantic Entropy (Farquhar et al., 2024) requires $S$ stochastic sequences per prompt at inference time (typically $S\in[5,20]$). The cost is dominated by the autoregressive decoding steps:
-$\mathcal{O}_{\text{SE}}\approx S\times(T_{\text{pre}}+T_{\text{gen}})\times C_{\text{LLM}}+\mathcal{O}_{\text{NLI}}(S^{2})$
-where $\mathcal{O}_{\text{NLI}}$ represents the quadratic cost of comparing all sample pairs using an entailment model.
-This linear scaling with $S$ introduces latency that is often unacceptable for interactive agents.
-Scalar Probes (PCP, SEP).
-Scalar probes operate on prompt representations extracted from a single forward pass of the base model.
-The probe itself is a lightweight classifier or regressor applied to a cached hidden state, without requiring additional sampling or decoding:
-$\mathcal{O}_{\text{SEP}}\approx 1\times T_{\text{pre}}\times C_{\text{LLM}}+C_{\text{MLP}}$
-While efficient, these methods predict a compressed scalar statistic and discard the geometric structure of the model’s semantic output distribution.
-Semantic Self-Distillation (SSD).
-SSD matches the inference complexity of scalar probes while retaining the distributional richness of sampling methods.
-The MDN head introduces a slightly larger parameter set than a linear probe, but this cost ($C_{\text{MDN}}$) is negligible compared to the transformer backbone.
-Crucially, the entropy calculation is analytical and does not require sampling:
-$\mathcal{O}_{\text{SSD}}\approx 1\times T_{\text{pre}}\times C_{\text{LLM}}+C_{\text{MDN}}$
-Thus, SSD effectively moves the computational burden from inference time (where latency matters) to training time, enabling $O(1)$ uncertainty estimation during deployment.
-Appendix D Ablations and Architecture Analysis
-D.1 PCA Dimension
+Appendix B Further Analyses
+In this section we provide additional analyses supporting observations made in the main text.
+Unless otherwise specified, results are reported on the TriviaQA dataset, where SSD exhibited the strongest performance variation across base models.
+B.1 Distillation Fidelity vs. Detection Performance
+We examine whether the quality of distillation (how closely the student learns the teacher distribution) constrains hallucination detection performance.
+We define distillation fidelity ($\rho_{\text{fidelity}}$) as the Spearman rank correlation, over $N=1000$ test prompts, between the student’s predicted entropy and the teacher dispersion.
+A high $\rho_{\text{fidelity}}$ indicates a more faithful recovery of the teacher’s prompt-induced uncertainty ordering.
+Figure A1 plots hallucination detection AUROC against distillation fidelity across base models and student hyperparameter settings.
+We observe a strong positive relationship, with a cross-model Spearman correlation of $\rho_{\text{meta}}=0.65$, indicating that detection performance is limited by the student’s ability to learn the prompt-to-dispersion mapping.
+The failure cases, notably Qwen3-4B, exhibit near-zero $\rho_{\text{fidelity}}$, suggesting that the relevant uncertainty signal is not accessible from the chosen prompt representation.
+Figure A1: Distillation fidelity drives detection performance.
+The X-axis shows distillation fidelity ($\rho_{\text{fidelity}}$): the Spearman correlation between the student’s predicted entropy and the teacher dispersion across TriviaQA test prompts.
+Multiple points per model correspond to different student hyperparameter settings.
+$\rho_{\text{meta}}$ is the Spearman correlation between the X and Y variables; it shows that models where the student fails to learn the distribution (low $\rho_{\text{fidelity}}$) result in poor detection.
+B.2 Semantic Consensus from the Predicted Mixture
+In Section 2.2, we noted that modelling an explicit predictive density enables latent-space sampling and computation of mixture statistics such as the mean.
+In language models, the mixture mean can serve as a proxy for semantic consensus.
+Whereas autoregressive decoding produces a single trajectory through output space, the mixture mean aggregates across the model’s sampled distribution in embedding space.
+We evaluate whether this analytic mean more faithfully approximates the empirical semantic centroid of sampled answers than a single low-temperature decoding.
+We treat the empirical mean of $S=32$ stochastic samples in embedding space as a reference semantic centroid, then measure the mean squared distance (MSD) in Euclidean space between this centroid and both (i) the model’s low-temperature output and (ii) the analytical mean of the student mixture, $\bm{\mu}_{ssd}=\sum_{k}\pi_{k}\bm{\mu}_{k}$.
+Lower MSD indicates better recovery of the model’s aggregate semantic belief.
+MSD is computed as $\mathbb{E}_{i}\left[\|\hat{\mathbf{z}}_{i}-\bm{\mu}^{\text{true}}_{i}\|_{2}^{2}\right]$ over examples $i$ in the subset.
+We report results for the best-performing student configuration (Ministral 8B), reflecting the strong architectural dependence of distillation fidelity.
+As shown in Table A1, the SSD mixture mean provides a closer approximation to the semantic centroid than a single default decoding pass.
+This effect is most pronounced for incorrect answers, where SSD improves centroid recovery in over 76% of test cases.
+This aligns with prior work showing that hallucinated outputs tend to deviate from semantic consensus [lin2024generating, phillips2025geometricuncertaintydetectingcorrecting].
+Together with the OOD results in Table 3, this result suggests that SSD exposes a coherent semantic belief state from which both prior (entropy) and posterior (likelihood and distance-to-consensus) reliability signals can be derived.
+Table A1: Semantic consensus approximation with Ministral 8B.
+We report the mean squared distance (MSD) of the default answer embedding and the SSD predicted mean to the true semantic centroid (average of $S=32$ samples).
+Values denote mean MSD, with the bootstrap standard deviation expressed as a percentage of the mean shown as a subscript.
+Subset Default MSD $\downarrow$ SSD MSD $\downarrow$ Imp. (%) $\uparrow$ SSD Win Rate (%) $\uparrow$
+All 0.0153.2 0.0081.9 +45.5 58.3
+Correct 0.0134.4 0.0092.3 +32.3 48.8
+Incorrect 0.0184.2 0.0073.4 +63.0 76.5
+Appendix C Ablations
+We analyze two design axes underlying SSD: the dimensionality of the semantic embedding space and the capacity of the student mixture model. All ablations are conducted using the TriviaQA dataset.
+C.1 PCA Dimension
 Standard sentence embeddings are high-dimensional (e.g., $d_{z}=768$).
@@
 We apply PCA to reduce the target space to $d_{pca}\in\{16,32,64,128\}$.
-This introduces a trade-off: higher dimensions preserve more semantic detail (increasing the theoretical upper bound of detection performance, measured by teacher dispersion), but make the regression task significantly harder for the student.
-As shown in Figure 3 and Table 6, while the detection performance of teacher dispersion improves monotonically with dimensionality, the student’s performance (SSD) peaks at $d_{pca}=16$.
+This introduces a trade-off between semantic resolution and distillation fidelity: higher dimensions preserve more semantic detail (raising the teacher dispersion baseline), but make the prompt-to-density regression task harder for the student to learn with fixed data.
+As shown in Figure A2 and Table A2, while the detection performance of teacher dispersion improves monotonically with dimensionality, the student’s performance (SSD) peaks at $d_{pca}=16$.
 Beyond this point, the student fails to effectively distil the increasingly sparse density, leading to a drop in detection accuracy.
@@
 For larger training datasets, we expect the SSD performance will continue to match or exceed the TD performance.
-Table 6: Effect of PCA dimensionality on Hallucination Detection (SSD) and OOD Verification. Scores are averaged across the best student configuration for each model.
+Table A2: Effect of PCA dimensionality on Hallucination prediction (SSD) and OOD Verification. Scores are averaged across the best student configuration for each model.
 PCA Dim Teacher Dispersion (TD) Peak SSD AUROC Peak OOD AUROC
@@
 128 0.757 $\pm$ 0.025 0.703 $\pm$ 0.029 0.950 $\pm$ 0.055
-Figure 3: As PCA dimension increases, the semantic resolution improves and the teacher dispersion baseline rises. The distillation task however becomes harder, causing SSD performance to fall.
-D.2 Student Capacity and Information Saturation
+Figure A2: As PCA dimension increases, the semantic resolution improves and the teacher dispersion baseline rises. The distillation task however becomes harder, causing SSD performance to fall.
+C.2 Student Capacity
 We investigate whether the performance gap on difficult models can be closed by increasing the student’s capacity. We analyze two axes of complexity:
@@
     Backbone Capacity: We vary the MLP hidden dimension $H\in\{128,256,512,1024\}$ and depth $D\in\{2,3,4,6\}$.
-Table 5 reveals that the optimal hyperparameters vary by base model.
-Aggregating results across all models (Table 7) reveals capacity saturation, where increasing the network width or depth yields negligible gains in mean AUROC.
-This saturation suggests that the bottleneck is not the student’s expressivity, but the mutual information available in the final-token prompt representation.
-The student appears to have extracted most of the readily accessible signal; further improvements would require richer inputs (e.g., attention-pooled representations) rather than larger students.
+Table A4 reveals that the optimal hyperparameters vary by base model.
+Aggregating results across all models (Table A3) reveals capacity saturation, where increasing the network width or depth yields negligible gains in mean AUROC.
+This saturation indicates that performance is not limited by student expressivity, but by the information available in the prompt representation used for distillation.
+The student appears to extract most of the accessible uncertainty signal from this representation; further improvements would require richer inputs (e.g., attention-pooled representations) rather than larger students.
 Although we observe optimal performance for $K=10$ mixture components, we opt not to increase this further given our set number of $S=32$ samples; modelling more than ten unique semantic concepts in this setting is unrealistic and would likely damage the generalization ability of the trained student.
-Table 7: Marginal impact of student capacity (Fixed PCA=16).
+Table A3: Marginal impact of student capacity (Fixed PCA=16).
 Scores represent the mean AUROC across all runs with the specified parameter; standard deviations indicate variability across different base models and remaining hyperparameter configurations.
@@
 6 0.656 $\pm$ 0.048
-Figure 4: Student capacity ablation, averaged across all seven models investigated.
-Appendix E Full Hallucination Detection Results
-Table 8: Hallucination Detection AUROC.
+Figure A3: Student capacity ablation on the TriviaQA dataset, averaged across all seven models investigated.
+Appendix D Model Configuration and Selection
+Here we summarize model characteristics and specific configuration choices used throughout our experiments.
+Table A4 reports base model accuracy on TriviaQA, the optimal probe layers identified for PCP and SEP, and the student architecture that achieved the highest SSD AUROC for each model.
+Table A4: Model-specific configuration details on TriviaQA.
+‘Accuracy’ denotes base model QA accuracy.
+‘Total’ gives the number of transformer layers, while ‘PCP’ and ‘SEP’ report the selected probe layers.
+‘Best SSD Config’ lists the student architecture (HH=hidden size, DD=depth, KK=mixture components) achieving the highest SSD AUROC.
+Layers Best SSD Config
+Model Accuracy Total PCP SEP H D K AUROC
+Qwen3-8B 0.620 36 32 24 128 6 10 0.677
+Qwen3-4B 0.565 36 18 20 256 2 10 0.699
+Llama 3.1 8B 0.700 32 20 16 256 4 2 0.687
+Llama 3.2 3B 0.585 28 15 19 512 3 5 0.737
+Ministral 8B 0.656 36 33 34 1024 6 1 0.748
+SmolLM3 3B 0.567 36 20 19 512 6 5 0.710
+Gemma 3 4B 0.532 34 22 19 512 2 10 0.700
+D.1 Layer Selection Protocol
+To select the student input layer $\ell^{*}$, we follow kossen2024semanticentropyprobesrobust and train linear probes on all layers to predict binarized Semantic Entropy.
+We choose the layer that maximizes validation accuracy on this proxy task.
+This heuristic assumes that layers most predictive of scalar dispersion also contain the richest signal for learning the full semantic density.
+In practice, the selected layers tend to lie in deeper parts of the network (Table A4).
+Appendix E Computational Complexity Analysis
+The primary motivation for Semantic Self-Distillation is to enable generic uncertainty estimation in AR models without sampling.
+Here we compare the asymptotic complexity of our method against sampling-based baselines.
+Let $T_{\text{pre}}$ be the prompt length, $T_{\text{gen}}$ be the generation length, $S$ be the number of samples, and $C_{\text{LLM}}$ be the cost of a single transformer forward pass per token.
+Sampling-based Methods (SE, TD).
+Computing Semantic Entropy [farquhar_detecting_2024] requires $S$ stochastic sequences per prompt at inference time (typically $S\in[5,20]$).
+The cost is dominated by the autoregressive decoding steps:
+$\mathcal{O}_{\text{SE}}\approx S\times(T_{\text{pre}}+T_{\text{gen}})\times C_{\text{LLM}}+\mathcal{O}_{\text{NLI}}(S^{2})$
+where $\mathcal{O}_{\text{NLI}}$ represents the quadratic cost of comparing all sample pairs using an entailment model.
+This linear scaling with $S$ introduces latency that is unacceptable in many applications, for example in interactive agentic systems.
+Scalar Probes (PCP, SEP).
+Scalar probes operate on prompt representations extracted from a single forward pass of the base model.
+The probe itself is a lightweight classifier or regressor applied to a cached hidden state, without requiring additional sampling or decoding:
+$\mathcal{O}_{\text{SEP}}\approx 1\times T_{\text{pre}}\times C_{\text{LLM}}+C_{\text{MLP}}$
+While efficient, these methods predict a compressed scalar statistic and discard the geometric structure of the model’s semantic output distribution.
+Semantic Self-Distillation (SSD).
+SSD matches the inference complexity of scalar probes while retaining the distributional richness of sampling methods.
+The MDN head introduces a slightly larger parameter set than a linear probe, but this cost ($C_{\text{MDN}}$) is negligible compared to the transformer backbone.
+Crucially, the entropy calculation is analytical and does not require sampling:
+$\mathcal{O}_{\text{SSD}}\approx 1\times T_{\text{pre}}\times C_{\text{LLM}}+C_{\text{MDN}}$
+Thus, SSD effectively moves the computational burden from inference time (where latency matters) to training time, enabling $O(1)$ uncertainty estimation during deployment.
+Appendix F Full Hallucination Prediction Results
+We report full per-model hallucination detection results corresponding to the summaries in Tables 2 and 1. These tables provide AUROC and AUPRC metrics for all methods across datasets.
+Table A5: Hallucination Prediction AUROC.
 Values report mean AUROC (%) over 1,000 bootstrap resamples of the test set, with the bootstrap standard deviation shown as a subscript.
-We mark the best overall result in bold and the best single-pass result with an underline.
-SSD achieves competitive performance with teacher dispersion but lags scalar probes and SE.
-Model PCP SEP SSD (Ours) TD SE
-Qwen3-8B 76.91.5 76.21.5 67.71.8 68.41.7 81.61.4
-Qwen3-4B 80.21.4 78.31.5 69.91.6 73.21.6 78.51.5
-Llama 3.1 8B 77.01.6 74.51.7 68.62.0 69.51.7 85.21.3
-Llama 3.2 3B 76.71.5 75.71.5 73.81.7 66.51.7 80.71.4
-Ministral 8B 74.01.7 77.11.5 74.81.6 74.31.6 80.51.5
-SmolLM3 3B 79.71.4 78.91.4 71.11.6 69.81.6 83.21.2
-Gemma 3 4B 75.41.5 70.71.7 70.11.7 67.61.7 73.11.6
-Table 9: Hallucination Detection AUPRC.
+We mark the best method per dataset and model in bold.
+TriviaQA MMLU
+Model PCP SEP SE TD SSD PCP SEP SE TD SSD
+Qwen3 8B 76.91.5 76.21.5 81.61.4 68.31.6 67.71.8 69.81.6 59.51.8 63.71.6 63.51.7 57.11.8
+Qwen3 4B 80.21.4 78.31.4 78.61.4 73.21.6 69.91.6 67.91.7 64.31.8 60.51.6 63.11.8 64.91.7
+Llama 3.1 8B 77.01.7 74.41.7 85.11.3 69.61.7 68.71.9 71.51.6 64.91.7 70.61.6 69.01.6 63.81.7
+Llama 3.2 3B 76.71.5 75.71.6 80.61.4 66.61.7 73.71.6 66.01.9 64.41.8 70.11.7 67.11.8 63.81.7
+Ministral 8B 74.01.7 77.11.6 80.51.5 74.31.7 74.91.6 67.01.7 62.81.8 69.31.7 65.11.8 63.61.8
+SmolLM3 3B 79.61.4 78.71.4 83.11.2 69.91.6 71.11.6 67.01.7 66.31.8 68.81.7 62.31.8 62.01.8
+Gemma 3 4B 75.21.6 70.61.6 73.11.5 67.61.6 70.01.7 66.21.8 57.81.8 61.31.6 62.81.7 60.71.8
+Table A6: Hallucination Prediction AUPRC.
 Values report mean AUPRC (%) over 1,000 bootstrap resamples of the test set, with the bootstrap standard deviation shown as a subscript.
-We mark the best overall result in bold and the best single-pass result with an underline.
-SSD consistently outperforms the teacher dispersion (TD) baseline, effectively shifting the cost of dispersion estimation from inference time to training time.
-Model PCP SEP SSD (Ours) TD SE
-Qwen3-8B 64.92.7 64.92.7 57.02.8 52.52.7 75.02.1
-Qwen3-4B 74.32.3 70.82.5 61.82.5 59.62.5 73.32.1
-Llama 3.1 8B 59.63.1 54.03.0 51.93.0 44.32.8 73.52.7
-Llama 3.2 3B 68.92.4 65.42.6 65.72.7 55.82.7 74.22.4
-Ministral 8B 59.22.9 62.92.8 59.92.9 57.92.9 69.62.7
-SmolLM3 3B 75.12.2 73.32.2 62.32.6 56.92.5 78.91.9
-Gemma 3 4B 69.42.4 65.02.5 66.22.5 63.22.4 69.22.3
+We mark the best method per dataset and model in bold.
+TriviaQA MMLU
+Model PCP SEP SE TD SSD PCP SEP SE TD SSD
+Qwen3 8B 64.82.6 64.92.7 75.02.1 52.42.6 56.92.8 66.92.3 56.32.3 60.32.2 61.92.3 55.92.3
+Qwen3 4B 74.32.4 70.82.5 73.42.2 59.72.4 61.92.5 60.12.5 57.62.5 53.42.3 58.52.4 59.42.5
+Llama 3.1 8B 59.73.1 54.03.0 73.42.7 44.42.7 51.92.9 72.72.1 62.62.3 70.52.1 70.92.1 62.72.4
+Llama 3.2 3B 68.92.3 65.42.6 74.12.2 55.82.6 65.52.5 76.41.9 74.52.0 78.41.8 74.02.0 74.51.9
+Ministral 8B 59.23.0 62.92.8 69.52.8 57.92.8 59.92.8 72.92.2 70.72.1 76.01.9 71.42.1 71.72.1
+SmolLM3 3B 75.02.1 73.12.2 78.81.9 57.12.5 62.42.5 67.72.3 67.62.2 70.52.2 64.62.2 63.02.3
+Gemma 3 4B 69.22.5 64.72.5 69.22.2 63.12.3 66.12.5 73.52.1 63.82.1 67.61.9 71.12.0 69.22.1